### v2

# Assortment Optimization Under History-Dependent Effects

[…]
Assortment Optimization Under History-Dependent Effects
[…]
This paper examines how to plan multi-period assortments when customer utility depends on historical assortments. We formulate this problem as a nonlinear integer programming problem and propose solution methodologies for different regimes of this problem. First, we present a sequential revenue-ordered policy and show that it is optimal when historical assortments positively affect customer utility. Second, we show that the problem is NP-hard under the presence of a negative history-dependent effect (such as a satiation effect). In this case, we propose using a lifting-based framework to reformulate the problem as a mixed-integer exponential cone program that state-of-the-art solvers can solve. Additionally, we identify an optimal cyclic policy for an asymptotic regime, and we also relate its length to the customer’s memory length. Finally, we present a case study using a catering service dataset that shows that our model demonstrates good fitness and can effectively balance variety and revenue.
[…]
1 Introduction
Many companies sell products to customers who make repeat purchases or visit over multiple periods. For example, corporate cafeterias provide workday luncheons for employees, gyms offer various classes to their members every day, and in live streaming sales, streamers plan product assortments for consecutive broadcasting sessions to fans. Many studies have explored and validated that customers’ memory of historical experiences may affect their willingness to buy or retain a product (Bowden 2009, Lemon and Verhoef 2016). For example, a customer may initially feel satisfied with a product, but this satisfaction may wane over time. On the other hand, a customer may also prefer to consistently consume the same product, which leads to sustained or increased satisfaction. How to incorporate these historical effects and offer various period-by-period product assortments is therefore a critical decision for assortment planning.
Studies have shown that consideration of customers’ history-dependent effects in pricing decisions can significantly improve revenue (e.g., Wang 2016, Chen et al. 2023c). However, history-dependent assortment planning problems are underexplored due to data scarcity and the decision complexity. Many managers continue to rely only on their experience, industry knowledge, and a rough estimate of the demand in order to manually determine assortments, and ignore the impact of historical assortments on customer behavior. With increasing collection of transactional data, companies seek to leverage these data to understand and estimate customers’ preferences to build data-driven assortment planning systems. For instance, new technology companies have emerged that revolutionize traditional cafeterias by establishing membership systems, gathering and analyzing transactional data, and offering data-driven customized lunch menus on a daily basis, which leads to increased revenue and customer engagement. For further details and references in this regard, the reader is referred to Fooda’s Cafeteria Solution [^1] https://www.fooda.com/cafeteria-solutions-corporate-dining and Zerocater’s Corporate Cafeteria Solution. [^2] https://zerocater.com/corporate-cafeterias/
Planning assortments using history-dependent effects is challenging because decision makers need to account for the interplay between cross-product and cross-period effects. Cross-product effects refer to the influence of other products in an assortment on purchase decisions. For instance, customer substitution behavior is commonly modelled using discrete choice models, like the multinomial logit (MNL) model (Gallego et al. 2019). Cross-period effects refer to the temporal influences on customer utility, i.e., the utility of a product in a given period depending on its offering history. Given both types of effect, the offering decisions for multiple products over multiple periods are made simultaneously. The number of candidate assortments, therefore, scales exponentially with the number of products and the planning horizon. This can be enormous for even a practically-sized problem. Moreover, modelling customer behavior under both types of effect makes the decision more complex because, for a given period, the purchase probability is a nonlinear function—specifically, a softmax function—of the history-dependent utility functions.
Figure 1: The benefits of incorporating history-dependent effects into assortment planning
Despite the complexity of assortment planning with history-dependent effects, the benefit of considering such effects is significant. Using the data from a corporate cafeteria with which we cooperated, we evaluate assortments derived from the MNL model, from a history-dependent model (our model [^3] We estimate the history-dependent effect and memory length (i.e., two periods) using transactional data of a corporate cafeteria. Please refer to Section 5 for details.), and the current practice observed in the data (as a benchmark to normalize the revenue). The results shown in Figure 1 indicate that our model achieves the highest average daily revenue in out-of-sample test, approximately 19% more than the revenue generated by the MNL model. Our model can also produce assortments that have greater variety than those produced by the MNL model. The current practice decided by managers resulted in assortments with a greater variety but also the lowest revenue. See Section 5 for detailed discussion on the case study.
1.1 Description of Results
We study a multi-period assortment planning problem in which a firm facing both cross-product and cross-period effects needs to decide what to offer in each period. We model customer choice in a given period using a variant of the MNL model in which the utility of each product linearly depends on its past offering decisions. In our model, an assortment decision for a given period affects both current and future purchase behavior in a nonlinear fashion. In other words, our assortment planning problem is a nonlinear integer programming problem.
We propose solution methodologies for different regimes of this problem. First, we present a sequential revenue-ordered (RO) policy and show that it is optimal if historical assortments positively affect customer utility, i.e., there is an addiction effect. This characterization indicates that our problem with the addiction effect is solvable in polynomial time. Second, we prove that when historical assortments negatively affect the current utility, i.e., there is a satiation effect, the problem becomes NP-hard even when there are only two planning horizons. In this case, we propose a mixed-integer exponential cone program (MIECP) formulation to find a global optimal solution. The key idea behind our formulation is to lift the original nonconvex constraints into a higher-dimensional space where the lifted constraints are simpler to model. This simpler reformulation allows us to utilize prevalent techniques in the optimization literature, such as the McCormick envelope, the perspective formulation, and the Lovász extension, to model the lifted structures. Numerical studies indicate that our formulation efficiently achieves optimality and yields higher revenue than heuristic policies, especially under strong satiation effects. For problems with a long-term planning horizon, theoretically, we show that a cyclic policy is optimal, and find that, under a certain non-overlapping condition, the optimal cycle length equals to the customer memory length plus one. Computationally, finding such cyclic policies is NP-hard but also admits an MIECP formulation. Moreover, for the case with a memory length of one, we derive a mixed-integer linear programming (MILP) formulation that is stronger than the conic formulation.
We conduct a case study using transactional data from a corporation cafeteria operated by a high-tech catering service company in China. The cafeteria primarily serves workday lunches for employees who work nearby. We estimate customers’ utility functions and then design weekly day-by-day lunch menus. The estimation shows that a satiation effect does exist, but it only lasts for a short period (i.e., two or three days). We demonstrate that our history-dependent choice model can provide menus that balance variety and revenue, while the menus used in practice sacrifice revenue and the menus proposed by the classic MNL model have low variety and low revenue.
We highlight the contributions of our paper as follows:
  • 1.
    Our model, which we refer to as HAP, simultaneously incorporates the cross-period and cross-product effects into the multi-period assortment planning process. We demonstrate via a case study of a corporate cafeteria that the operations in our model have several benefits. First, our result suggests that the service provider should provide a greater variety of products if satiation effects exist. When satiation effects last longer, it is beneficial to offer more types of products across periods. Furthermore, it is recommended that a product be offered only once within the memory length in order to prevent a utility decrease due to satiation effects, particularly in the case of products with high prices. This observation naturally leads to a “non-overlapping” cyclic assortment planning policy, which we explore in Section 4, where we establish its optimality. Most importantly, our history-dependent policy effectively balances variety and the revenue objective, resulting in assortments that achieve the highest revenue in out-of-sample tests while maintaining the same variety as found in practical scenarios.
  • 2.
    We characterize, in terms of history-dependent effects, the conditions under which a sequential revenue-ordered policy globally solves our HAP model. In general, it can be shown that finding a global optimal solution is NP-hard. To tackle this difficulty, we develop an MIECP formulation of our HAP model by using lifting and coupling ideas from the optimization literature, including perspective formulation, the McCormick envelope, and the Lovász extension. Computational studies show that our formulation and its variant outperform a MILP formulation of (HAP) obtained using a prevalent strategy in the mixed-integer nonlinear programming (MINLP) literature and the original nonlinear integer formulation (HAP) solved with the state-of-the-art MINLP solver SCIP.
  • 3.
    We establish the asymptotical optimality of cyclic policies, and finding such policies allows an MIECP formulation. Our study reveals a connection between the length of an optimal cyclic policy and the length of customer memory. In particular, we characterize the condition under which the optimal cyclic length is the customer memory length plus one. This condition, in addition, leads to an MILP formulation for finding an optimal two cyclic policy, which is, theoretically, tighter than the previous conic formulation.
1.2 Literature Review
Extensive research has incorporated history-dependent effects into operations problems, including pricing and promotion (Wang 2016, Bi et al. 2020, Chen et al. 2023c), service design (Bernstein et al. 2022, Lei et al. 2023, Fan et al. 2024), resource allocation (Adelman and Mersereau 2013), retention management (Bastani et al. 2022, Kanoria et al. 2024), etc. The research shows that ignoring history-dependent effects in customer behavior may lead to substantial losses. For instance, considering a reference price in choice estimation can reduce a $5.82\%$ prediction error (Wang 2018), and the optimal bundle accounting for customer satiation obtains a revenue lift greater than $4.5\%$ (Gürlek et al. 2023).
History-dependent effects are receiving increasing attention in assortment planning. For instance, Chen et al. (2023b) study personalized assortment planning problems when customers’ choice probabilities depend on historical purchases. Several recent studies focus on designing online assortment algorithms to learn customers’ satiation levels and maintain high engagement (e.g., Cao et al. 2019, Bastani et al. 2022, Sumida and Zhou 2023). Our study investigates the choice-based offline assortment planning problem, such as the daily menu design for corporate cafeterias. Unlike online recommendation settings, the offline assortment planning problem determines multi-period assortments at the beginning of a planning horizon.
A prevalent strategy for solving assortment optimization is to construct revenue-ordered assortments. Talluri and Van Ryzin (2004) show that the optimal solution of the assortment optimization problem under the MNL model is revenue-ordered. Wang and Wang (2017) propose a quasi-revenue-ordered policy to solve assortment optimization with endogenous network effects. Gao et al. (2021) prove that the revenue-ordered structure is optimal for the assortment optimization problem under the MNL model when the offered assortment is gradually revealed. Xu and Wang (2023) exploit the revenue-ordered property to solve a multistage assortment optimization problem. We contribute to this literature by showing that a sequential revenue-ordered policy is optimal for our problem under the addiction effect.
[…]
Last, we comment that our paper contributes to a growing set of applications of conic programming/discrete optimization in designing operations systems, including joint inventory-location problems (Atamtürk et al. 2012), designing battery swap networks for electric vehicles (Mak et al. 2013), scheduling (Kong et al. 2013), process flexibility design (Yan et al. 2018, Simchi-Levi and Wei 2012), electric vehicle charging (Chen et al. 2023a), proactive policing (He et al. 2023a), network flow (Simchi-Levi et al. 2019).
1.3 Structure
The rest of the paper is organized as follows. In Section 2, we formally define the assortment planning problem, show its NP-hardness, and identify a case that is solvable in polynomial time. In Section 3, we propose an equivalent MIECP formulation. In Section 4, we present our results on cyclic policies. We present a case study using real data in Section 5 and report numerical results on synthetic data in Section 6. We conclude the paper in Section 7. All proofs are given in the Appendix.
2 Model
We define our multi-period assortment planning problem with history-dependent effects in Section 2.1. In Section 2.2, we then show the NP-hardness of our HAP model when history-dependent effects are negative. In Section 2.3, we identify a case that is solvable in polynomial time. Specifically, we show in Theorem 2.2 that a sequential revenue-ordered policy is optimal if the history-dependent effects are positive.
2.1 Assortment Planning under History-Dependent Effects
Consider a firm that sells $N$ products to customers. Let $[N]:=\{1,\ldots,N\}$ [^4] In this paper, $[k]:=\{1,2,\dots,k\}$ for integer $k\in\operatorname{\mathbb{Z}}_{+}$. denote the set of all products, and let $[N]^{+}:=[N]\cup\{0\}$, where we use the index 00 to denote the outside option. For each product $i\in[N]$, let $r_{i}$ denote the revenue obtained from selling an instance of product $i$. We assume that products are indexed such that $r_{1}\geq r_{2}\geq\dots\geq r_{N}>0$. There is a total of $T$ planning horizons, and the firm needs to decide what to offer in each period. More specifically, for each period $t\in[T]$, let $\boldsymbol{x}^{t}=(x^{t}_{1},\ldots,x^{t}_{N})\in\{0,1\}^{N}$ be a binary decision variable modelling the subset of products offered in period $t$, i.e., $x^{t}_{i}=1$ if and only if product $i$ is offered in period $t$. We assume that in each period, one unit mass of customers arrives, and the customers select an item from the offered products or leave without a purchase. For example, in a setting such as corporate cafeterias, the firm serves relatively stable customers (e.g., most customers are employees who work nearby), and customers make purchase decisions on a daily basis.
To model the customer choice behavior in each period, we modify the MNL model to encapsulate the impact of the historical assortments on the current choice behavior. Under the random utility model (Gallego et al. 2019, Section 4), a customer assigns a random utility to each product and selects the product that yields the highest utility. When modeling the utility of each product, we take into account information about the history of assortments. More specifically, the utility of product $i$ in period $t$ affinely depends on its historical offering records within a memory length $M$:
$\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta_{i}^{m}+\xi_{i}^{t},$ (Utility)
where $M\in\operatorname{\mathbb{Z}}_{+}$ denotes the customer memory length and $\xi_{i}^{t}$ is a random variable. For simplicity in the presentation, we assume throughout this paper that $x_{i}^{k}:=0$ for $k\leq 0$. The parameter $\beta^{0}_{i}$ is the base utility of the product $i$ that can be determined by various factors such as brand, flavor, price, and so on. The parameter $\beta_{i}^{m}$ is the incremental utility if product $i$ is offered in the $m^{\text{th}}$ period before the current time. [^5] In our planning models, the assortments are fixed for each period. Therefore, we do not consider personalized assortments or model individual-level utility functions. Let $\boldsymbol{\beta}:=(\beta_{i}^{m})_{i\in[N],m\in[M]}$ denote the history-dependent effect matrix. Note that the value of $\beta_{i}^{m}$ can be either positive or negative. If $\beta_{i}^{m}$ is positive, exposure to product $i$ increases a customer’s preference for it and thus functions as a stimulus (Fox et al. 1997). Conversely, if $\beta_{i}^{m}$ is negative, product $i$ has a satiation effect, which leads to customer fatigue from repeated offerings.
As in the MNL model, we assume that random variables $\xi_{i}^{t}$ are i.i.d. as the standard Gumbel distribution. For a given assortment planning $\boldsymbol{x}:=(\boldsymbol{x}^{1},\cdots,\boldsymbol{x}^{T})$, we use $\pi^{t}_{i}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ to denote the probability of purchasing product $i$ in period $t$. Under the MNL model, the probability of purchasing product $i$ in period $t$ is
$\pi^{t}_{i}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ $=\dfrac{x^{t}_{i}\exp\bigl{(}\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta^{m}_{i}\bigr{)}}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}}\quad\operatorname{for}i\in[N],$
$\pi^{t}_{0}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ $=\dfrac{1}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}}.$
The purchase probabilities depend on two components: the current assortment $\boldsymbol{x}^{t}$ and the past assortments $\boldsymbol{x}^{t-1},\dots,\boldsymbol{x}^{t-M}$. The former dependence captures cross-product effects such as the substitution effect among products, while the latter dependence models cross-period effects.
With these definitions, assortment planning under history-dependent effects can be expressed as the following fractional binary program, called the history-dependent assortment planning problem (HAP):
$\max\Biggl{\{}\sum_{t\in[T]}\frac{1}{T}\sum_{i\in[N]}r_{i}\pi^{t}_{i}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})\Biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\,\operatorname{for}t\in[T]\text{ and }\boldsymbol{x}\in\mathcal{P}\Biggr{\}},$ (HAP)
where $\mathcal{X}$ (resp. $\mathcal{P}$) is a system of linear inequalities that models cross-product constraints (resp. cross-period constraints). The cross-product constraints often include resource restrictions within a single period, such as cardinality and capacity constraints, and address various operational considerations in the assortment planning literature (Sumida et al. 2021, Chen et al. 2024). Cross-period constraints describe the intertemporal relationships of assortments across different periods, such as non-overlapping assortments over consecutive periods, which naturally occur in multi-period assortment planning problems (Liu et al. 2020, Chen et al. 2023b).
2.2 Hardness Result
This subsection shows that our model (HAP) is NP-hard to solve when history-dependent effects are negative. In particular, when products have identical revenue and the planning horizon is two, we can reduce a NP-hard problem, the 3/4-Partition problem, to our problem (HAP).
Proposition 2.1
Our model (HAP) is NP-hard even when the planning horizon is two and the memory length is one.
2.3 Revenue-Ordered Policies and Optimality Structures
A prevalent strategy for solving assortment optimization is to construct RO assortments, which are a family of nested assortments selected according to their revenue-decreasing order. Specifically, RO structures and their variants have been shown to be optimal for solving assortment optimization problems under MNL choice models (Talluri and Van Ryzin 2004), sequential MNL choice models (Gao et al. 2021), and MNL choice models with network effects (Wang and Wang 2017). Inspired by the simplicity and optimality of RO structures, we would like to understand how it works in our problem.
We find that when there are no cross-product and cross-period constraints, if the history-dependent effect is non-negative, that is, $\boldsymbol{\beta}\geq 0$, the optimal assortment in each period is a RO assortment and the assortment planning can be obtained by a sequential revenue-ordered policy. The sequential RO policy is obtained by sequentially selecting the best RO assortment based on the currently updated product utility and maximizing revenue in the current period. This policy is adaptive while ignoring the impact of the current assortment decision on future revenue. Details of the sequential RO policy is shown in Algorithm 1.
Algorithm 1 Sequential revenue-ordered policy of (HAP)
1:  Input: $\mathbf{r},(\beta_{i}^{0})_{i\in[N]},\boldsymbol{\beta},M,N,T$
2:  for all $t=1,2,\dots,T$ do
3:     Update history assortments: $\boldsymbol{x}^{t-1},\dots,\boldsymbol{x}^{t-M}$, and set $\boldsymbol{x}^{t-m}=\boldsymbol{0}$ if $t-m\leq 0$.
4:     Compute the best revenue-ordered assortment $\boldsymbol{x}^{t}=\mathop{\mathrm{argmax}}\limits_{\boldsymbol{z}\in\{0,1\}^{N}}\sum_{i\in[N]}r_{i}\pi^{t}_{i}(\boldsymbol{z},\boldsymbol{x}^{t-1},\ldots,\boldsymbol{x}^{t-M})$
5:  end for
6:  Output: $\boldsymbol{x}^{t},t\in[T]$
Because there are no cross-product and cross-period constraints, the computation time required to find the optimal RO assortment is $\mathcal{O}(N)$ in each period (Talluri and Van Ryzin 2004, Liu and Van Ryzin 2008). Thus, the total computation time of the sequential RO policy is $\mathcal{O}(NT)$.
We are now in a position to state the main result of this subsection, Theorem 2.2, characterizing conditions on which the sequential RO policy is optimal.
Theorem 2.2
Assume that there are no cross-product or cross-period constraints. Then, the sequential-revenue-ordered policy solves model (HAP) if the history-dependent effects are non-negative, that is, $\boldsymbol{\beta}\geq\boldsymbol{0}$.
For the case of negative history-dependent effects, which is the computationally intractable regime of our model, as we argued in Proposition 2.1, the sequential policy fails to solve our model. However, the RO structure still exists and appears in a different form. For an optimal assortment planning of model (HAP), we argue that the set of products that are offered across all periods has a RO structure.
Proposition 2.3
Assume that there are no cross-product or cross-period constraints. Then, there exists an optimal assortment planning $\boldsymbol{x}=(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{T})$ of model (HAP) such that $\bigl{\{}i\in[N]\bigm{|}\sum_{t\in T}x^{t}_{i}\geq 1\bigr{\}}$ is revenue-ordered.
We conclude this section with an example to numerically illustrate that the optimality gap of RO-type policies can be significant for a general instance of (HAP). This motivates us to study the global optimality of (HAP) in Section 3.
Example 2.4
In this example, we compare the revenue of two heuristic policies and the best feasible solution obtained within 3600 seconds for randomly generated instances. The revenue and base utility of an product are randomly generated from uniform distributions with ranges of $[1,10]$ and $[-1,1]$, respectively. The history-dependent effect $\beta_{i}^{m}$ is uniformly sampled from $[-2,-1]$. One of the heuristic policies is the sequential RO-policy. The other is a repeated policy, a non-adaptive policy that chooses an optimal assortment for the first period and repeatedly offers the same assortment for the remaining $T-1$ periods. The relative revenue gap between a myopic policy and the feasible solution is
$G_{\text{rev}}=100\%\times\frac{R_{\mathsf{fea}}-R_{\mathsf{myopic}}}{R_{\mathsf{fea}}},$
where $R_{\mathsf{fea}}$ is the revenue generated by the best feasible solution of (HAP) achieved within 3600s and $R_{\mathsf{myopic}}$ is the revenue generated by one of two myopic policies.
Table 1: Revenue gap of myopic policies
[…]
N T $M=1$ $M=2$ $M=1$ $M=2$
30 5 8.26 14.86 13.84 35.66
30 10 10.21 26.12 15.78 46.41
50 5 7.17 12.48 12.93 32.71
50 10 27.18 22.12 32.34 42.35
70 5 5.14 7.85 10.71 23.05
70 10 6.37 14.45 12.03 29.71
[…]
3 Mixed-Integer Programming Formulation of (HAP)
In this section, we aim to solve the history-dependent assortment problem (HAP) to its global optimality. Due to the combinatorial and nonconvex nature of (HAP), state-of-the-art global optimization solvers such as BARON (Tawarmalani and Sahinidis 2005) and SCIP (Bestuzheva et al. 2021) fail to solve practical-sized instances—see Section 6.1 for computational evidence. Nevertheless, we focus on developing techniques for obtaining mixed-integer programming formulations of (HAP), which will allow us to leverage modern mixed-integer program solvers such as GUROBI and MOSEK. Then, in Section 4, we will use these techniques to find asymptotically optimal cyclic policies.
Our formulation introduces the following additional variables to represent the customer choice behavior. For each period $t\in[T]$, we use $\rho^{t}$ to denote the no-purchase probability for period $t$:
$\rho^{t}=\frac{1}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}},$
and for each product $i\in[N]$, we use $y^{t}_{i}$ to denote the choice probability for product $i$ in period $t$:
$y^{t}_{i}=\frac{x^{t}_{i}\exp\bigl{(}\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta^{m}_{i}\bigr{)}}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}}.$
In what follows, we refer to $\rho^{t}$ as the no-purchase probability variable and $y^{t}_{i}$ as the choice probability variable. Using these definitions, we can reformulate  (HAP) as follows:
$\max_{\boldsymbol{x},\boldsymbol{y},\rho}\quad$ $\sum_{i\in[N]}\frac{1}{T}\sum_{t\in[T]}r_{i}y_{i}^{t}$
[…]
$\rho^{t}+\sum_{i\in[N]}y_{i}^{t}=1\quad$ $\operatorname{for}\,t\in[T]$
$y^{t}_{i}=\rho^{t}x^{t}_{i}\exp\Bigl{(}\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta^{m}_{i}\Bigr{)}\quad$ $\operatorname{for}\,t\in[T]\text{ and }i\in[N].$ (Choice)
The objective function is a non-negative weighted combination of choice probability variables. The second constraint ensures that the sum of the choice probability and the no-purchase probability equals one for each period. The last constraint represents the choice probability for product $i$ in period $t$, and thus we refer to it as constraint  (Choice). This reformulation encapsulates the nonconvexity of problem  (HAP) in constraint (Choice). Hence, it suffices to derive a convex representation of constraint (Choice).
The main component of constraint (Choice) is the composition of a univariate exponential function and an affine utility function, defined as follows:
$\alpha_{i}(z^{1},\ldots,z^{M}):=\exp\biggl{(}\beta_{i}^{0}+\sum_{k\in[M]}\beta^{k}_{i}\cdot z^{k}\biggr{)}\qquad\operatorname{for}\;(z^{1},\ldots,z^{M})\in\{0,1\}^{M}.$
For a given assortment $\boldsymbol{x}:=(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{T})$, we interpret $\alpha_{i}(x^{t-1},\ldots,x_{i}^{t-M})$ as the attractiveness of product $i$ in period $t$. Therefore, we will henceforth refer to $\alpha_{i}(\cdot)$ as the attraction value function. To handle the discrete nature of the attraction value function, we will use the notion of a continuous extension. Given a function $f:\{0,1\}^{n}\to\mathbbm{R}$, a function $g:[0,1]^{n}\to\mathbbm{R}$ is called a continuous extension of $f(\cdot)$ if $g(\cdot)$ is continuous and $g(\boldsymbol{z})=f(\boldsymbol{z})$ for every $\boldsymbol{z}\in\{0,1\}^{n}$. In addition, if an extension function $g(\cdot)$ is convex (resp. concave), we say that it is a convex (resp. concave) extension. It is common for a discrete function to admit many convex (resp. concave) extensions. A natural idea is to use the tightest convex/concave extension. Since the discrete domain $\{0,1\}^{n}$ is the vertex set of the continuous domain $[0,1]^{n}$, the pointwise largest convex (resp. smallest concave) extension of $f(\cdot)$ is the convex (resp. concave) envelope, denoted as $\operatorname{conv}(f)(\cdot)$ (resp. $\operatorname{conc}(f)(\cdot)$), of $f(\cdot)$ over the box $[0,1]^{n}$ (see Tawarmalani and Sahinidis 2002, Theorem 6). Recall that the convex (resp. concave) envelope of $f(\cdot)$ is referred to the tightest under (resp. over)-estimator of $f$ over $[0,1]^{n}$ (Rockafellar 1997).
3.1 A Lifted Representation of (Choice)
In this subsection, we focus on modelling constraint (Choice). The idea is to lift the original nonconvex constraint (Choice) into a higher-dimensional space where the lifted constraint is dramatically simpler to represent than the original constraint. Specifically, this will allow us to invoke prevalent techniques such as the McCormick envelope (McCormick 1976), the perspective formulation (Günlük and Linderoth 2010), and the Lovász extension (Lovász 1983) to model the lifted nonconvex structures. We elaborate on this in the following Sections 3.1.1 and 3.1.2.
3.1.1 Lifting for Simplicity
[…]
$\lambda+\lambda f(\boldsymbol{x})\leq 1,$
where $\lambda\geq 0,\boldsymbol{x}\in\mathbbm{R}^{n}$, and $f(\cdot)$ is a convex function. This constraint can be relaxed into a convex constraint as follows. First, we multiply and divide $\boldsymbol{x}$ by $\lambda$ to obtain
$\lambda+\lambda f\Bigl{(}\frac{\lambda\cdot\boldsymbol{x}}{\lambda}\Bigr{)}\leq 1,$
where $\lambda f(\frac{\lambda\cdot\boldsymbol{x}}{\lambda})=0$ when $\lambda=0$. Next, we replace $\lambda\boldsymbol{x}$ with a new variable vector $\boldsymbol{u}$. This leads to
$\boldsymbol{u}=\lambda\boldsymbol{x}\quad\text{and}\quad\lambda+\lambda f\Bigl{(}\frac{\boldsymbol{u}}{\lambda}\Bigr{)}\leq 1,$
where the first constraint can be linearized using the McCormick envelope McCormick (1976) and the second constraint is convex, since the function $\lambda f(\boldsymbol{u}/\lambda)$ is the perspective function of the convex function $f(\cdot)$ (Rockafellar 1997). Formally, given a function $f:D\to\mathbbm{R}$, where $D$ is a bounded subset of $\mathbbm{R}^{n}$, the perspective function of $f$, denoted as $\operatorname{pers}(f)$, is defined as
$\operatorname{pers}(f)(\lambda,\boldsymbol{x})=\begin{cases}\lambda f(\frac{\boldsymbol{x}}{\lambda})&\text{ for }\lambda>0\text{ and }\boldsymbol{x}\in\lambda\cdot D\\ 0&\text{ for }\lambda=0\text{ and }\boldsymbol{x}=\boldsymbol{0}.\end{cases}$
Perspective reformulations have been widely used in solving nonconvex optimization problems (Günlük and Linderoth 2010); see Bestuzheva et al. (2023) for a review of applications and a detailed computational study of perspective reformulations.
We now present our lift of the constraint (Choice). Note that the choice probability of product $i$ in period $t$ is computed by scaling its attraction value function $\alpha_{i}(\cdot)$ with a nonlinear factor $\rho^{t}x_{i}^{t}$, that is,
$y^{t}_{i}=(\rho^{t}x^{t}_{i})\cdot\alpha_{i}(x_{i}^{t-1},\ldots,x_{i}^{t-M}).$
First, we introduce a new variable $\gamma_{i}^{t}$ to represent the nonlinear factor $\rho^{t}x_{i}^{t}$. Second, we use the new variable $\gamma_{i}^{t}$ to scale the historical assortment decisions $x^{t-m}_{i}$ with $m\in[M]$, and we introduce the variable $z_{im}^{t}$ to represent the scaled variable $\gamma_{i}^{t}x^{t-m}_{i}$. Using the additional variables $\gamma_{i}^{t}$ and $\boldsymbol{z}_{i}^{t}=(z_{i1}^{t},\ldots,z_{iM}^{t})$ and the definition of a perspective function, we decompose constraint (Choice) into constraints
$\gamma_{i}^{t}=\rho^{t}x_{i}^{t}$ (2a)
$z_{im}^{t}=\gamma_{i}^{t}x^{t-m}_{i}\quad\operatorname{for}\,m\in[M]$ (2b)
$y^{t}_{i}=\operatorname{pers}(\alpha_{i})(\gamma^{t}_{i},\boldsymbol{z}_{i}^{t}).$ (2c)
Constraint (2a) lifts the nonlinear scaling factor $\rho^{t}x_{i}^{t}$ to $\gamma_{i}^{t}$; constraint (2b) lifts the scaled historical assortment decision $\gamma_{i}^{t}x_{ik}^{t}$ to $z_{ik}^{t}$; and constraint (2c) perspectifies the attraction value function in the space of the newly introduced variables $\gamma_{i}^{t}$ and $\boldsymbol{z}_{i}^{t}$.
To obtain a mixed-integer convex formulation of constraint  (Choice) in the lifted space, we individually relax each of the three constraints in (2). Specifically, we use the McCormick envelope (McCormick 1976) to relax constraint (2a). To apply the McCormick envelope to our setting, we derive a lower bound $\rho^{t}_{L}$ and upper bound $\rho^{t}_{U}$ on the no-purchase probability variable $\rho^{t}$, and impose the following constraints:
$\gamma_{i}^{t}$ $\leq\min\bigl{\{}\rho^{t}_{L}x_{i}^{t}+\rho^{t}-\rho^{t}_{L},\ \rho^{t}_{U}x_{i}^{t}\bigr{\}}$ (3)
$\gamma_{i}^{t}$ $\geq\max\bigl{\{}\rho^{t}_{L}x_{i}^{t},\ \rho^{t}_{U}x_{i}^{t}+\rho^{t}-\rho^{t}_{U}\bigr{\}}.$
Similarly, using 00 (resp. $\rho^{t}_{U}$) as a lower (resp. upper) bound on $\gamma^{t}_{i}$, the McCormick envelope reformulates constraint (2b) as follows:
$z_{im}^{t}$ $\leq\min\bigl{\{}\gamma_{i}^{t},\rho^{t}_{U}x^{(t-m)}_{i}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M]$ (4)
$z_{im}^{t}$ $\geq\max\bigl{\{}0,\ \rho^{t}_{U}x^{(t-m)}_{i}+\gamma^{t}_{i}-\rho^{t}_{U}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M].$
Finally, we represent constraint (2c) by using two continuous extensions of the attraction value function $\alpha_{i}(\cdot)$. On one side, a natural continuous extension $\tilde{\alpha}_{i}:[0,1]^{M}\to\mathbbm{R}$ can be obtained from $\alpha_{i}(\cdot)$ by relaxing the binary assortment variables to continuous ones. This continuous function is convex since it is the composition of a convex function and an affine function. Using this extension, we relax constraint (2c) to
$y^{t}_{i}\geq\operatorname{pers}(\tilde{\alpha}_{i})(\gamma^{t}_{i},\boldsymbol{z}^{t}_{i}).$
This is a convex inequality since $\tilde{\alpha}_{i}(\cdot)$ is a convex function and so is its perspective function (Rockafellar 1997). More precisely, this constraint can be represented as an exponential cone constraint (Ben-Tal and Nemirovski 2001) (see more details on obtaining an exponential cone representation in Appendix 9.3):
$y_{i}^{t}\geq\exp(\beta^{0}_{i})\cdot\gamma^{t}_{i}\cdot\exp\biggl{(}\frac{\sum_{m\in[M]}\beta_{i}^{m}z_{im}^{t}}{\gamma^{t}_{i}}\biggr{)},$ (5)
where the right-hand-side function can be regarded as a convex under-estimator of the choice probability variable $y^{t}_{i}$. On the other side, an exact representation of (2c) requires us to overestimate the choice probability variable. To do this, we replace $\alpha_{i}(\cdot)$ with its tightest concave extension or, equivalently, concave envelope. Using the concave envelope, we arrive at another convex relaxation of (2c):
$y_{i}^{t}\leq\operatorname{pers}\bigl{(}\operatorname{conc}(\alpha_{i})\bigr{)}(\gamma^{t}_{i},\boldsymbol{z}^{t}_{i}).$ (6)
In the following Section 3.1.2, we show that the concave envelope of $\alpha_{i}(\cdot)$ can be described using the Lovász extension of $\alpha_{i}(\cdot)$ (Lovász 1983).
3.1.2 The Lovász Extension of $\alpha_{i}(\cdot)$
We now derive a closed-form expression of the concave envelope of the attraction value function, which appears in constraint (6). To do this, we exploit the fact that the attraction value function $\alpha_{i}(\cdot)$ is supermodular. Recall that a function $f:\{0,1\}^{n}\to\mathbbm{R}$ is supermodular if
$f(x^{\prime}\vee x^{\prime\prime})+f(x^{\prime}\wedge x^{\prime\prime})\geq f(x^{\prime})+f(x^{\prime\prime})\quad\operatorname{for}x^{\prime},x^{\prime\prime}\in\{0,1\}^{n},$
where $x^{\prime}\vee x^{\prime\prime}$ (resp. $x^{\prime}\wedge x^{\prime\prime}$) is the coordinate-wise maximum (resp. minimum) of $x^{\prime}$ and $x^{\prime\prime}$.
Lemma 3.1
If $\boldsymbol{\beta}\leq 0$ then the attraction value function $\alpha_{i}(\cdot)$ is supermodular over $\{0,1\}^{M}$.
Function $\alpha_{i}(\cdot)$ maps the offering decisions for product $i$ over the past $M$ periods to an attraction value. Given that the history-dependent effects are nonpositive, the more number of times that product $i$ has been offered in the past, the lower its current attraction value. Consequently, increasing the number of past offerings results in smaller changes to the attraction value due to the convexity of the attraction value function. This means that, as product $i$ is offered more in the past, the decay in its attraction value diminishes with each additional offering, indicating a complementary effect (Topkis 1998).
The supermodularity of $\alpha_{i}(\cdot)$ allows us to use the Lovász extension to describe its concave envelope (Lovász 1983, Tawarmalani et al. 2013, Atamtürk and Narayanan 2022). The Lovász extension is an extension of a function $f$ defined on $\{0,1\}^{n}$ to a function $\hat{f}$ defined on $[0,1]^{n}$. For each subset $S$ of $[n]$, let $\chi^{S}\in\{0,1\}^{M}$ be its indicator vector—that is, the $i^{\text{th}}$ coordinate of $\chi^{S}$ is $1$ if and only if $i\in S$. Observe that every vector $\boldsymbol{x}\in[0,1]^{n}$ can be expressed uniquely as $\boldsymbol{x}=\lambda_{1}\chi^{T_{1}}+\lambda_{2}\chi^{T_{2}}+\cdots\lambda_{k}\chi^{T_{k}}$, where $\lambda_{i}\geq 0$ for $i=1,\ldots,k$ and $\sum_{k=0}^{n}\lambda_{k}=1$, and $\emptyset\subseteq T_{0}\subset T_{1}\subset\cdots\subset T_{k}\subseteq[n]$. Thus,
$\hat{f}(\boldsymbol{x})=\lambda_{1}f(\chi^{T_{1}})+\lambda_{2}f(\chi^{T_{2}})+\cdots+\lambda_{k}f(\chi^{T_{k}})$
is a well-defined extension of the function $f$ (called the Lovász extension of $f$) on the continuous domain $[0,1]^{n}$. The Lovász extension serves as a bridge between supermodular functions and concave functions: $f$ is supermodular if and only if its Lovász extension $\hat{f}$ is concave (Lovász 1983). Moreover, when the Lovász extension $\hat{f}$ is concave, it is expressible as the pointwise minimum of affine functions (Tawarmalani et al. 2013). In the following lemma, we show that when these results are scaled with the variable $\gamma^{t}_{i}$, we can obtain an explicit polyhedral description of constraint (6).
Lemma 3.2
Let $\Omega$ be the set of permutations of memory length $\{1,2,\ldots,M\}$. For a permutation $\sigma\in\Omega$, define a collection of nested historical offerings in $\{0,1\}^{M}$ as follows: $\boldsymbol{h}^{\sigma}_{0}=(0,0,\ldots,0)$ and $\boldsymbol{h}^{\sigma}_{k}=\boldsymbol{h}^{\sigma}_{k-1}+\boldsymbol{e}_{\sigma(k)}$ for $k\in[M]$, where $\boldsymbol{e}_{k}$ is the $k^{\text{th}}$ principal vector in $\mathbbm{R}^{M}$. Then, under the presence of constraint (4), constraint (6) is equivalent to the following system of linear inequalities:
$y^{t}_{i}\leq\alpha_{i}(\boldsymbol{h}^{\sigma}_{0})\cdot\gamma^{t}_{i}+\sum_{k\in[M]}\bigl{(}\alpha_{i}(\boldsymbol{h}^{\sigma}_{k})-\alpha_{i}(\boldsymbol{h}^{\sigma}_{k-1})\bigr{)}\cdot z^{t}_{ik}\qquad\operatorname{for}\sigma\in\Omega.$ (7)
3.2 The Final Formulation
Using the linear constraints (3), (4), and (7) and the exponential cone constraint (5) to represent constraint (Choice), we obtain an MIECP formulation of problem (HAP) as follows:
$\max_{\boldsymbol{x},\boldsymbol{y},\rho}$ $\sum_{i\in[N]}\frac{1}{T}\sum_{t\in[T]}r_{i}y_{i}^{t}$ (Conic)
[…]
$\rho^{t}+\sum_{i\in[N]}y_{i}^{t}=1\quad$ $\operatorname{for}\,t\in[T]$
[…]
Theorem 3.3
(Conic) is a mixed-integer exponential cone formulation of problem (HAP).
Remark 3.4
In this remark, we consider the case where the memory length is less than two. In this case, we are able to derive the tightest convex extension, also known as the convex envelope, of $\alpha_{i}(\cdot)$. Using the convex envelope instead of the natural continuous extension, we obtain the following tighter formulation than the one given by constraint (5):
$y_{i}^{t}\geq$ $\bigl{(}\alpha_{i}(\boldsymbol{e}_{1})-\alpha_{i}(\boldsymbol{0})\bigr{)}\cdot z_{i1}^{t}+\bigl{(}\alpha_{i}(\boldsymbol{e}_{2})-\alpha_{i}(\boldsymbol{0})\bigr{)}\cdot z_{i2}^{t}+\alpha_{i}(\boldsymbol{0})\gamma_{i}^{t}$ (8a)
$y_{i}^{t}\geq$ $\bigl{(}\alpha_{i}(\boldsymbol{1})-\alpha_{i}(\boldsymbol{e}_{2})\bigr{)}\cdot z_{i1}^{t}+\bigl{(}\alpha_{i}(\boldsymbol{1})-\alpha_{i}(\boldsymbol{e}_{1})\bigr{)}\cdot z_{i2}^{t}+\bigl{(}\alpha_{i}(\boldsymbol{e}_{1})+\alpha_{i}(\boldsymbol{e}_{2})-\alpha_{i}(\boldsymbol{1})\bigr{)}\cdot\gamma_{i}^{t},$ (8b)
[…]
We conclude this section with a discussion of how our formulation relates to existing formulations. Recently, a large amount of research has focused on developing strong formulations for binary linear fractional programming problems and MNL-based assortment planning problems (Sen et al. 2018, Mehmanchi et al. 2019, Atamtürk and Gómez 2020, Kılınç-Karzan et al. 2023, He et al. 2023b, Chen et al. 2024). To apply these results in solving problem (HAP), we need to linearize the nonlinear functions appearing in denominators and numerators—specifically, the attraction value functions. A prevalent linearization technique replaces each attraction value function by its multilinear extension, called the Fourier Expansion in O’Donnell (2014), and then linearizes the resulting multilinear functions using the recursive McCormick relaxation (McCormick 1976, Khajavirad 2023). We combine this linearization strategy with binary linear fractional programming formulation techniques to obtain an exact formulation of problem (HAP), which we refer to as the multilinear-extension-based formulation. The complete derivation is provided in Appendix 11.1. Later, in Section 6.1, a computational experiment shows that our formulation (Conic) significantly outperforms the multilinear-extension-based formulation.
[…]
$\max_{\boldsymbol{a}\in\mathcal{A}\subseteq\{0,1\}^{n}}\quad\sum_{k=1}^{K}\lambda_{k}\cdot\frac{\exp(u_{k}(\boldsymbol{a}))}{1+\exp(u_{k}(\boldsymbol{a}))},$
[…]
4 Cyclic Policies: Optimality and Computation
In this section, we consider the situation in which a firm plans assortments for a long period. We establish the asymptotic optimality of cyclic policies in Section 4.1 and formulate the cyclic policy problem as an MIECP in Section 4.2. In Section 4.3, we characterize the conditions under which the length of optimal cyclic policies is the memory length plus one. Finally, in Section 4.4, we show that such conditions lead to an MILP formulation for finding an optimal two-cyclic policy that is theoretically tighter than the conic formulation.
4.1 Optimality of Cyclic Policies
The assortment planning problem with an infinite horizon can be formulated as follows. The firm can select any assortment sequence (or policy) $\boldsymbol{x}=\{\boldsymbol{x}^{t}\}_{t\in\operatorname{\mathbb{Z}}_{+}}$ with elements in $\{0,1\}^{N}$. As before, the firm may face cross-period and cross-product constraints. Let $\mathcal{P}^{\infty}$ denote the feasible set of cross-period constraints (e.g., a product cannot be offered in two consecutive periods), and let $\mathcal{X}$ denote the feasible set of cross-product constraints (e.g., assortment cardinality constraints). Then the firm aims to find an optimal assortment sequence for the following problem:
$\sup_{\boldsymbol{x}}\biggl{\{}\lim_{T\to\infty}\frac{1}{T}\sum_{t\in[T]}\sum_{i\in[N]}r_{i}\pi_{i}^{t}(\boldsymbol{x}^{t},\dots,x^{t-M})\biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\;\operatorname{for}t\in\operatorname{\mathbb{Z}}_{+},\ \boldsymbol{x}\in\mathcal{P}^{\infty}\biggr{\}}.$ (Infty)
An assortment policy $\boldsymbol{x}=\{\boldsymbol{x}^{t}\}_{t\in\operatorname{\mathbb{Z}}_{+}}$ is cyclic if there exists a positive integer $L$ such that $\boldsymbol{x}^{t+L}=\boldsymbol{x}^{t}$ for all $t\in\operatorname{\mathbb{Z}}_{+}$. The smallest $L>0$ for which this holds is called the cycle length of $\boldsymbol{x}$. A cyclic policy $\boldsymbol{x}$ can be represented by a finite sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$, where $L$ is the cyclic length of $\boldsymbol{x}$. In contrast, whenever we say that a finite sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$ is cyclic, we are referring to the policy in which this finite sequence of assortments is repeated infinitely often.
For a given cycle of length $L$, we are interested in finding a cyclic policy of length $L$ that maximizes the long-term average revenue. This is equivalent to searching for a sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$ of length $L$ that maximizes the average revenue by repeating the sequence. Such a sequence can be defined as an optimal solution of a variant of (HAP). To define the variant, we modify the attraction value function so that it depends on past assortments in the cycle. Given a position $t$ in the cycle of length $L$, we use a function $\tau(\cdot|t)$ to track positions within the memory length $M$:
$\tau(m|t)=\begin{cases}{(t-m\text{ mod }L)}&\text{ if }(t-m\text{mod }L)>0\\ {L}&\text{ if }(t-m\text{ mod }L)=0.\end{cases}$
The attraction value of product $i$ at position $t$, which depends on the assortments of product $i$ offered over the memory length, can be specified as follows:
$\alpha_{i}^{t}(x^{1}_{i},\ldots,x^{L}_{i})=\exp\biggl{(}\beta_{i}^{0}+\sum_{m\in[M]}\beta^{m}_{i}x^{\tau(m\mid t)}_{i}\biggr{)}.$
A cyclic policy is called an $L$-cyclic policy if it is optimal for the following variant of problem (HAP):
$\max\Biggl{\{}\sum_{i\in[N]}\frac{1}{L}\sum_{t\in[L]}\frac{r_{i}x^{t}_{i}\alpha^{t}_{i}(x^{1}_{i},\ldots,x^{L}_{i})}{1+\sum_{j\in[N]}x^{t}_{j}\alpha^{t}_{j}(x^{1}_{i},\ldots,x^{L}_{i})}\Biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\operatorname{for}t\in[L],\ \boldsymbol{x}\in\mathcal{P}^{L}\Biggr{\}},$ (Cycle)
where $\mathcal{X}$ denotes cross-product constraints and $\mathcal{P}^{L}$ denotes cross-period constraints. In addition, we assume that a finite sequence of assortments satisfying $\mathcal{P}^{L}$ can be repeated infinitely often to obtain a cyclic policy that satisfies the cross-period constraint $\mathcal{P}^{\infty}$ in (Infty).
The main result of this subsection relates problems  (Infty) and (Cycle). To establish this result, we represent (Infty) as a directed graph, which we refer to as the assortment graph. Each node of the graph denotes a list of $M$ assortments, indexed from the latest to the earliest assortments in memory. Each arc represents a memory transition driven by assortment decisions. Specifically, an arc starting at a node should end at a node that drops the earliest assortment of the starting node and adds a new assortment. Hence, each arc encodes an assortment decision in a period and its weight is defined as the single-period revenue generated by the new assortment. With such a graph, we can show our assortment planning problem (Infty) is equivalent to finding the maximum mean cycle of the assortment graph (Karp 1978), whose cycle length is denoted as $L^{*}$. Given $L^{*}$, we can use (Cycle) to compute the cyclic policy. We formalize this observation as the following theorem.
Theorem 4.1
Given an instance of (Infty), there exists a positive integer $L^{*}$ such that $L^{*}$-cyclic policy is optimal to (Infty).
A detailed proof is provided in the appendix. Formulation techniques developed for solving (HAP) can be used to obtain an MIECP formulation for (Cycle); see details in Section 4.2
4.2 Finding Optimal Cyclic Policies via a Mixed-Integer Exponential Cone Program
Formulation techniques developed for solving (HAP) can be used to obtain a mixed-integer exponential cone formulation for (Cycle). Since the main discrepancy between two models is the definition of attraction value functions, we only need to modify constraints in the formulation (Conic) that involve attraction value functions. In particular, we replace $\alpha_{i}(\cdot)$ in constraint (5) and (7) by $\alpha^{t}_{i}(\cdot)$, obtaining
$y^{t}_{i}$ $\geq\exp(\beta^{0}_{i})\cdot\gamma^{t}_{i}\cdot\exp\Biggl{(}\frac{\sum_{m\in[M]}\beta_{i}^{m}z_{im}^{\tau(m\mid t)}}{\gamma^{t}_{i}}\Biggr{)}$ (9)
$y^{t}_{i}$ $\leq\alpha^{t}_{i}(\boldsymbol{h}^{\pi}_{0})\cdot\gamma^{t}_{i}+\sum_{m\in[M]}\bigl{(}\alpha^{t}_{i}(\boldsymbol{h}^{\pi}_{m})-\alpha^{t}_{i}(\boldsymbol{h}^{\pi}_{m-1})\bigr{)}\cdot z_{im}^{\tau(m\mid t)}\qquad\operatorname{for}\pi\in\Pi.$
This replacement yields the following mixed-integer exponential cone formulation of (Cycle):
$\max_{\boldsymbol{x},\boldsymbol{y},\rho}$ $\sum_{i\in[N]}\frac{1}{L}\sum_{t\in[L]}r_{i}y_{i}^{t}$ (Cycle-Conic)
$\operatorname{s.t.}$ $\boldsymbol{x}\in\mathcal{P}^{L}\text{ and }\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}$ $\operatorname{for}t\in[L]$
$\rho^{t}+\sum_{i\in[N]}y_{i}^{t}=1$ $\operatorname{for}\,t\in[L]$
[…]
(Cycle-Conic) has its own value because it is an efficient way to generate a feasible assortment planning with a large planning horizon. Especially, the retailer may pre-determine a cycle length $L$ based on experience or industry knowledge and use (Cycle-Conic) to derive the optimal policy under a given $L$.
4.3 Cycle Length Characterization Under the Non-overlapping Condition
It is intuitive that, in order to avoid the negative impact on a product’s utility if the product exhibits a satiation effect, a product should not be offered consecutively. In particular, we say that a cross-period constraint on a policy $\boldsymbol{x}$ is non-overlapping if $M+1$ adjacent assortments $(\boldsymbol{x}^{t},\boldsymbol{x}^{t+1},\dots,\boldsymbol{x}^{t+M})$ are non-overlapping, that is, for each product $i\in[N]$, $x^{t}_{i}+x^{t+1}_{i}+\dots+x^{t+M}_{i}\leq 1$ for every $t\in\operatorname{\mathbb{Z}}_{+}$. This non-overlapping condition is common in sequential assortment decisions (Liu et al. 2020, Chen et al. 2023b). Li et al. (2024) identify a similar non-overlapping condition on customer orders to design a polynomial algorithm for the e-commerce assortment planning problem. Under the non-overlapping condition, any cyclic policy with length $L\leq M$ must have empty sets in some periods. We do not consider such cyclic policies. For cyclic policies with length $L>M$, we find that this condition allows us to relate the length of an optimal cycle policy for solving (Infty) to the customers’ memory length as follows.
Theorem 4.2
Assume that $M<N$. An $(M+1)$-cycle policy is optimal for problem (Infty) if the cross-period constraint is non-overlapping.
The condition $M<N$ guarantees that there at least exists a non-overlapping assortment planning with each assortment non-empty. Without the non-overlapping condition, the length of the optimal cyclic policy varies as the parameters change. In particular, a slight fluctuation in the input parameters, like the base utility of a product, could change the length of the optimal cycle in the absence of the non-overlapping condition. In Appendix 10.4, we provide an example to illustrate this situation.
The base utility and history-dependent effect are estimated from real data and lie in confidence intervals. Hence, given the fact that the optimal cycle length is sensitive to input parameters, finding a cycle length with robust performance may be more critical than characterizing the optimal cycle length in practice. A positive result is that although the $(M+1)$-cyclic policy is not necessarily optimal for problem  (Infty) without the non-overlapping constraint, it still provides a good approximation for solving our problem, as illustrated in the following example.
Example 4.3
In this example, we numerically test the performance of the $(M+1)$-cyclic policy. We fix the number of products $N=30$ and vary the length of the planning horizon $T\in\{10,30,50\}$. The revenue and base utility of each product are randomly generated from uniform distributions with ranges $[1,10]$ and $[-1,1]$, respectively. For a weak (W) (resp. strong (S)) satiation effect, we set the history-dependent effect $\beta_{i}^{m}$ using a uniform distribution $U[-1,0]$ (resp. $U[-2,-1]$). For each instance, we compute the revenue gap between the $(M+1)$-cyclic policy and the best feasible solution, defined as follows:
$G_{\textsf{rev,cycle}}=100\%\times\frac{R_{\mathsf{fea}}-R_{\mathsf{cycle}}}{R_{\mathsf{fea}}},$
where $R_{\mathsf{cycle}}$ is the revenue achieved by the $(M+1)$-cyclic policy and $R_{\mathsf{fea}}$ is the revenue of the best feasible solution obtained by (Conic) within $7200$ seconds.
Table 2: Revenue gap of cyclic policies
[…]
W 10 0.51 1.69 S 10 0.39 4.45
W 30 0.16 0.60 S 30 -0.06 -0.50
W 50 0.08 0.52 S 50 -0.52 0.46
[…]
The non-overlapping condition provides an additional benefit in problem formulation when cycle length is small. In the following subsection, we provide a bound-free model formulation when $M=1$ with non-overlapping requirement.
4.4 A Strong Formulation for Finding Two-Cyclic Policies
Here, we will focus on the computation of the two-cyclic policy with the non-overlapping constraint. Although the formulation (Cycle-Conic) is convenient, one can directly solve it to obtain an optimal cyclic policy. However, its computational performance depends on the bounds of no-purchase probabilities. We next propose a bound-free formulation for the two-cyclic policy, which is provably tighter than its conic counterpart (Cycle-Conic), and, in Section 6.2, we demonstrate its computational efficacy. Our formulation is built on the polyhedral structure of the relation between choice probabilities across two periods and assortment decisions.
We start by reformulating (Cycle) into a simpler form which is a special case of the assortment optimization problem studied in Caro et al. (2014). When $M=1$, attraction value functions are univariate functions over the binary domain $\{0,1\}$, and, thus, can be represented using an affine function. For each product $i\in[N]$, let $u_{i}:=\exp(\beta^{0}_{i})$ and $d_{i}:=\exp(\beta^{0}_{i}+\beta^{1}_{i})-\exp(\beta^{0}_{i})$. Then, for each period $t\in[T]$, the attraction value function can be linearized as follows:
$\alpha^{1}_{i}(x^{2}_{i})=u_{i}+d_{i}x^{2}_{i}\quad\text{ and }\quad\alpha^{2}_{i}(x^{1}_{i})=u_{i}+d_{i}x^{1}_{i}\quad\operatorname{for}\,(x^{1}_{i},x^{2}_{i})\in\{0,1\}^{2},$
where $u_{i}$ represents the base attraction value of product $i$, and $d_{i}$ is the incremental attraction value of product $i$ conditional on whether it is offered in another period. Due to the presence of the non-overlapping constraint $x_{i}^{1}+x_{i}^{2}\leq 1$, we obtain a further simplification:
$x_{i}^{t}\alpha_{i}^{t}(x_{i}^{\tau(1|t)})=u_{i}x_{i}^{t}+d_{i}x_{i}^{1}x_{i}^{2}=u_{i}x_{i}^{t}\qquad\text{ for }(x^{1}_{i},x^{2}_{i})\in\{0,1\}^{2},$
where the second equality holds since $x^{1}_{i}x_{i}^{2}=0$. Therefore, (Cycle) is equivalent to
$\max$ $\sum_{t=1,2}\frac{1}{2}\sum_{i\in[N]}\frac{r_{i}u_{i}x^{t}_{i}}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}$ (2-Cyclic)
$\operatorname{s.t.}$ $x_{i}^{1}+x_{i}^{2}\leq 1$ $\operatorname{for}i\in[N]$
$\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}$ $\operatorname{for}t=1,2.$
Note that this reformulation reveals a connection between (Cycle) and the assortment problem studied in Caro et al. (2014) in which a retailer must decide, in advance, the release date of each product in a given collection over a selling season. More specifically, (2-Cyclic) is exactly the case where a retailer faces two planning horizons and sells products with a single period life-cycle, and, in each period, customers make a purchase decision on offered products according to the MNL model.
Next, we introduce variables for building a strong formulation of (2-Cyclic). For each period $t=1,2$, we use $\rho^{t}$ and $\gamma^{t}_{i}$ to represent the no-purchase probability and choice probability of product $i$, respectively
$\rho^{t}=\frac{1}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}\quad\text{ and }\quad\gamma_{i}^{t}=\frac{x^{t}_{i}}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}.\qquad\operatorname{for}i\in[N]$ (CC)
Note that this is known as the Charnes-Cooper transformation (Charnes and Cooper 1962). For each period $t=1,2$, it maps all possible assortments $\boldsymbol{x}^{t}\in\{0,1\}^{N}$ into all possible choice probabilities $(\rho^{t},u_{1}\gamma^{t}_{1},\ldots,u_{N}\gamma^{t}_{N})$, where $(\rho^{t},\boldsymbol{\gamma}^{t})$ belongs to
$\rho^{t}+\sum_{i\in[N]}u_{i}\gamma^{t}_{i}=1\quad\text{ and }\quad\gamma^{t}_{i}\in\{0,\rho^{t}\}\,\operatorname{for}i\in[N].$ (10)
Using these relations, we are ready to present a new formulation of (2-Cyclic):
$\max\quad$ $\sum_{t=1,2}\frac{1}{2}\sum_{i\in[N]}r_{i}u_{i}\gamma_{i}^{t}$
$\operatorname{s.t.}\quad$ $x^{1}_{i}+x^{2}_{i}\leq 1\text{ and }\boldsymbol{x}^{t}\in\mathcal{X}\{0,1\}^{N}$ $\operatorname{for}i\in[N]\text{ and }t=1,2$ (11a)
$\rho^{t}+\sum_{i\in[N]}u_{i}\gamma^{t}_{i}=1\text{ and }0\leq\gamma^{t}_{i}\leq\rho^{t}\quad$ $\operatorname{for}i\in[N]\text{ and }t=1,2$ (11b)
$x^{t}_{i}=\gamma^{t}_{i}+\sum_{j\in[N]}u_{j}\Gamma^{t}_{ij}$ $\operatorname{for}i\in[N]\text{ and }t=1,2$ (11c)
$\max\{0,\gamma^{t}_{i}+\gamma^{t}_{j}-\rho^{t}\}\leq\Gamma^{t}_{ij}\leq\min\{\gamma^{t}_{i},\gamma^{t}_{j}\}\quad$ $\operatorname{for}i\in[N]\,,j\in[N]\setminus\{i\},\text{ and }t=1,2$ (11d)
$\Gamma^{t}_{ii}=\gamma^{t}_{i}$ $\operatorname{for}i\in[N]\text{ and }t=1,2.$ (11e)
The constraint (11b) is obtained from relaxing the constraints from (10). The constraint (11c) is derived from the following equality, with the variable $\gamma^{t}_{i}$ replacing the first term and a new variable $\Gamma^{t}_{ij}$ replacing each fractional term in the summation,
$x^{t}_{i}=\frac{x^{t}_{i}}{1+\sum_{k\in[N]}u_{k}x^{t}_{k}}+\sum_{j\in[N]}u_{j}\frac{x_{i}^{t}x_{j}^{t}}{1+\sum_{k\in[N]}u_{k}x^{t}_{k}}\qquad\operatorname{for}\,t=1,2.$
To obtain the constraint (11d), we use the no-purchasee variable $\rho^{t}$ to scale the following McCormick envelope of $x_{i}^{t}x_{j}^{t}$ (McCormick 1976):
$\max\{0,x^{t}_{i}+x^{t}_{j}-1\}\leq x^{t}_{i}x^{t}_{j}\leq\min\{x_{i}^{t},x_{j}^{t}\}\qquad\operatorname{for}j\in[N].$
The last constraint (11e) is obtained by observing that the relation $x^{t}_{i}x^{t}_{i}=x^{t}_{i}$ holds for binary variables. Unlike the formulation (Cycle-Conic), this formulation does not rely on bounds of the no-purchasing probability variables. More specifically, in the derivation of (Cycle-Conic), we use $\rho^{t}x^{t}_{i}$ as a nonlinear scaling factor, demanding a linearization using bounds of $\rho^{t}$. In the development of linear inequalities in (11), we avoid using nonlinear scaling factor and, instead, use the no-purchasing probability variables, $(\rho^{1},\rho^{2})$, as scaling variables to derive (11d).
The following result shows this bounds-free formulation is theoretically tighter than the formulation (Cycle-Conic). One of the most important properties of a MIP formulation is the strength of its natural continuous relaxation that is obtained by ignoring the integrality constraints. This is important because a tighter continuous relaxation often indicates a faster convergence of the branch-and-bound algorithm, on which most commercial MIP solvers are built (Vielma 2015). Since (HAP) has a maximization objective, a tighter formulation has a smaller continuous relaxation objective value. Let $Z_{\textsc{Bounds-free}}$ (resp. $Z_{\textsc{Cycle-Conic}}$) be the optimal objective value of the natural continuous relaxation of formulation (11) (resp. (Cycle-Conic)). The two objective values have the following relationship.
Theorem 4.4
For $M=1$, $Z_{\textsc{Bounds-free}}\leq Z_{\textsc{Cycle-Conic}}$ if the cross-period constraint is non-overlapping.
5 Case Study
This section presents a case study using transaction data from a corporation cafeteria. We first estimate the history-dependent effect using the historical data and then use the results to design assortments. Our estimations indicate that customers have a significant but short-term satiation effect. Furthermore, our findings suggest that including history-dependent effects in choice models will lead to greater assortment varieties and higher revenue.
5.1 Data Description and Estimation
We cooperated with a Chinese company that provides catering services to over $4,000$ corporation cafeterias and were given access to a dataset for one cafeteria that mainly serves workday luncheons for customers who work nearby. [^6] The cafeteria is closed on holidays and weekends. The dataset includes $27,144$ transactions for $53$ main dishes, covering $110$ workdays from July to December, 2021. The data for each transaction include the time, dish, and quantity, as well as a card ID if the dish was purchased with a membership card. The dataset also includes information about the price, main ingredient, and flavor of each dish. In what follows, we use the terms dish and product and also the terms menu and assortment interchangeably.
Our dataset indicates that 84.8% of the transactions came from 418 members, which generated 84.5% of the total revenue. Thus, we focus on the members’ transactions because they are frequent visitors and generate most of the revenue. Our dataset also reveals that the probability of a member dining in the cafeteria more than once for two, three, or four consecutive days was 72.5%, 87.7%, and 91.2%, respectively. This suggests that customers’ dish selections could be influenced by their past dining experiences.
We predict customers’ choices using the MNL model with history-dependent effects. The utility function for product $i$ on day $t$ is
$U_{i}^{t}=\boldsymbol{\theta}^{T}X_{i}^{t}+\theta_{r}r_{i}+\sum_{m\in[M]}x_{i}^{t-m}\beta^{m}+\xi_{i}^{t},$
where $X_{i}^{t}$ denotes covariates that include the main ingredient and flavor of product $i$, product fixed effects, and month fixed effects; $r_{i}\in[0,1]$ is a normalized price; $\boldsymbol{\theta}$ and $\theta_{r}$ are coefficients of $X_{i}^{t}$ and price, respectively; $\boldsymbol{\theta}^{T}X_{i}^{t}+\theta_{r}r_{i}$ is the base utility $\beta_{i}^{0}$ defined in Section 2; and the $\xi_{i}^{t}$s are i.i.d. standard Gumbel random variables. The history-dependent effect is captured by the term $\sum_{m\in[M]}x_{i}^{t-m}\beta^{m}$, in which $M$ is the memory length. Our model is a standard MNL model if $M=0$. Note that we set the same history-dependent coefficient for all products because not all products have consecutive offering records and therefore we cannot estimate the product-wise history-dependent effect. We do not include the cross-product history-dependent effect because the estimation result shows that the cross-product effect is approximately one-tenth of the direct effect and is significant only when $M=1$. This is in line with numerous empirical findings that the direct-product effect is much more significant than the cross-product effect (e.g., Vilcassim et al. 1999, Dubé and Manchanda 2005, Mehta 2007). Therefore, we focus on the direct history-dependent effect.
We utilize members’ transactional data to calibrate the MNL choice model with the utility function presented above. We split the data into two sets: transaction data from July to September and transaction data from October to December. We first use five-fold cross-validation on the first data set to estimate the above utility function, and then we optimize weekly menus based on the estimators. Next, we conduct the five-fold cross-validation again, this time using the second data set, to select the value of $M$ with the best out-of-sample performance, and we set this as the “ground truth” model for evaluating revenue, which is a common approach in policy evaluation (Cao et al. 2022). To address the fact that our dataset lacks no-purchase records, we employ the expectation-maximization (EM) approach, which is a widely used technique to deal with censored data (Talluri and Van Ryzin 2006, van Ryzin and Vulcano 2017, Şimşek and Topaloglu 2018), to calibrate our model.
Table 3 summarizes the estimation and five-fold validation results for $M=\{0,1,2,3,4\}$, using the third quarter data. Column $\theta_{r}$ presents price coefficient estimators, and columns $\beta^{1}$ to $\beta^{4}$ give history-dependent effects. The columns labelled $\chi^{2}$ and Log-likelihood provide the average out-of-sample value in the five-fold cross-validation (normalized by sample size). The percentage values in parentheses indicate the relative improvement of each model with a history-dependent effect compared to the standard MNL model.
Table 3: Estimators and cross-validation results
Estimators Out-of-sample Performance
$M$ $\theta_{r}$ $\beta^{1}$ $\beta^{2}$ $\beta^{3}$ $\beta^{4}$ $\chi^{2}$ Log-likelihood
0 $-2.493^{***}$ 0.0916 -1.7985
1 $-2.468^{***}$ $-0.480^{***}$ 0.0900 ($1.75\%$) -1.7966 ($0.10\%$)
(0.082) (0.041)
2 $-2.444^{***}$ $-0.494^{***}$ $-0.322^{***}$ 0.0890 ($2.84\%$) -1.7953 ($0.18\%$)
(0.082) (0.041) (0.040)
3 $-2.440^{***}$ $-0.494^{***}$ $-0.307^{***}$ $-0.225^{***}$ 0.0896 ($2.18\%$) -1.7951 ($0.19\%$)
(0.082) (0.041) (0.041) (0.052)
4 $-2.433^{***}$ $-0.499^{***}$ $-0.307^{***}$ $-0.222^{***}$ $-0.106^{***}$ 0.0905 ($1.20\%$) -1.7955 ($0.17\%$)
(0.082) (0.041) (0.041) (0.052) (0.045)
  • Note: The numbers in parentheses under the estimators report the standard errors of the estimations. “-” indicates that a model does not include the corresponding predictor. ${***}p<0.001$. The columns “$\chi^{2}$” and “Log-likelihood” show the average value of the five-fold cross-validation (normalized by sample size) followed by a percentage in parentheses, which is the relative improvement of each model with a history-dependent effect compared to the standard MNL model.
Table 3 shows significant negative history-dependent effects, and the scale of the effects decrease as the time interval between the current and historical offerings increases. The estimation results are consistent for different values of $M$. For instance, $\beta^{1}$ is consistently about $-0.49$ for all $M$. This suggests that the attraction value of a product will shrink by about $\exp(-0.49)=61\%$ if it was offered yesterday. In other words, if a product is sold on two consecutive days, we need to lower its price by at least RMB $(0.49/2.493)\times 15\approx 2.95$ to compensate for the negative history-dependent effect. [^7] We normalize prices by the maximum price RMB $15$ in the estimation. These results demonstrate the importance of considering history-dependent effects in assortment planning.
The average out-of-sample validation results are reported in the last two columns of Table 3. They suggest that history-dependent effects can improve prediction power (i.e., lower $\chi^{2}$s and higher log-likelihoods). When $M=2$, the model presents the lowest $\chi^{2}$ and the second highest log-likelihood. This result suggests that a small memory length ($M=2$) can capture the history-dependent effect in customers’ choice.
5.2 Assortment Planning
In this section, we use the formulation proposed in Section 3 to optimize weekly menus using the above estimators. We design weekly workday menus because the cafeteria does not frequently operate on weekends, and in practice, the cafeteria indeed plans for a week. We randomly select a month [^8] Our estimation includes monthly fixed effects. Hence, the product’s utility will vary month by month. and use the estimated parameters in Table 3 as the input for our conic models. To keep the menu sizes comparable, we impose a daily cardinality constraint of $6$, which is the median assortment size in the dataset.
Figure 2: Offering times, price, and base attraction value of offered dishes
5.2.1 Assortment Patterns
Figure 2 shows the total number of products offered, the prices, and the base attraction values obtained for the assortments under the MNL model, under history-dependent MNL models with different memory lengths, and under an assortment planning used in current practice. The current practice assortment planning was randomly chosen from data on the same month which was used for computing the base attraction value when optimization (the base utility contains month fixed effects).
As shown in Figure 2 (a) and (b), the MNL-based model tends to offer fewer products with higher prices. Moreover, the number of products that are offered increases with the memory length $M$, because it is beneficial to offer more products and avoid repeating the same product within the memory length when satiation effects last for a long period. When the memory length increases, the assortments include more products with lower base attraction values, as shown in Figure 2 (c).
Figure 3: Assortment details
Figure 3 presents assortment details for examining how the memory length affects assortment patterns. Figure 3 (a)-(d) show the assortments obtained from MNL-based models with $M=0,1,2,3$. The horizontal axes represent products indexed from the highest to the lowest price. The vertical axes represent the days of a week. The products offered are colored blue. The subfigures show that the interval between two adjacent offerings of the same dish grows when $M$ increases, and all optimal assortments represent a cyclic-style structure. The interval for two adjacent offerings of most products is just the memory length $M$, which suggests that the cycle length is $M+1$, aligning with the $(M+1)$-cyclic policy in Section 4. Moreover, it can be observed that products with higher prices or lower base attraction values are more sensitive to the satiation effect, and thus will be offered at most once within the span of the memory length. For instance, the high-price products 1 to 4 strictly repeat in intervals of $M$ days to avoid the satiation effect, possibly because the revenue loss caused by the utility deterioration will be scaled up by the high price. On the other hand, products 6 and 12 are provided on two consecutive days when $M=2$ due to their relatively high base attraction value. The price and base attraction value for each product are shown in Figure 3 (e) and (f) for reference.
5.2.2 Performance Comparison
In this section, we focus on two metrics to evaluate the assortment performance: variety and revenue, as these are concerns for the corporation cafeteria. We adopt the Herfindahl-Hirschman Index (HHI), a measure of market concentration (Rhoades 1993), to quantify the variety of an assortment planning. The HHI value is $\sum_{i\in[n]}(\frac{h_{i}}{H})^{2}$, where $n$ is the total number of products, $h_{i}$ is the number of times product $i$ is offered, and $H=\sum_{i\in[n]}h_{i}$. A low HHI value reflects a large variety. To compute the out-of-sample revenue, we set the model with the best prediction performance for the second data set, i.e., the data set in the fourth quarter, as the “ground truth” model. Besides the assortments discussed in Section 5.2.1, we additionally compute assortments obtained by the $(M+1)$-cyclic policy with non-overlapping constraints that is presented in Section 4.
Figure 4: Assortment variety and revenue
Figure 4 illustrates the relationship between the variety and revenue of different assortments. The horizontal axis represents the HHI value, and the vertical axis represents the normalized out-of-sample average daily revenue. We normalize the out-of-sample revenue of different assortment plannings with the revenue of current practice as a benchmark. Generally, assortments with a low HHI value can help avoid the negative impacts of satiation effects and maintain high product utility. Figure 4 shows that the assortments generated by our model ($M>0$) have a lower HHI and higher revenue than the assortments obtained with the MNL model. For instance, the average revenue obtained using our model with $M=2$ is approximately $19\%$ greater than the revenue under the MNL model. The $(M+1)$-cyclic policy also performs well. With the same memory length $M$, the $(M+1)$-cyclic policy increases the assortment variety via the non-overlapping constraint and even generates more revenue than the assortments without non-overlapping constraints. We also note that, in practice, assortments that are determined manually have low HHI values but carry substantial revenue losses.
In conclusion, our case study first calibrates the history-dependent effect using transaction data from a corporation cafeteria. Then we demonstrate that offering a wide variety of products with a low frequency can efficiently compensate for the satiation effect. Finally, our findings highlight the advantages of our model, which captures the complex trade-off between products’ utility, their revenue, and history-dependent effects.
6 Numerical Study
This section presents the results of our numerical studies on synthetic data. In Section 6.1, we illustrate the efficiency of formulation (Conic) in terms of computation time and solution quality. In Section 6.2, we demonstrate the performance of cyclic policies. All our numerical studies were implemented in the Julia, version 1.8.2 (Bezanson et al. 2017), using modeling language JuMP (Lubin et al. 2023), on a Virtual Machine with 64GB RAM and a 32-core Intel Xeon (Skylake, IBRS) @2.39 GHz processor. All mixed-integer linear programs were solved using Gurobi 11.0.0 (Gurobi Optimization, LLC 2023) and mixed-integer exponentical cone programs were solved using Mosek 10.2 (MOSEK 2022).
6.1 Formulation Efficiency
We consider four alternate formulations for solving (HAP):
  • $\mathsf{Conic}$: This is the mixed-integer exponential cone formulation given by (Conic).
  • $\mathsf{Env}$: This is the mixed-integer linear formulation obtained by replacing the exponential cone constraint in (Conic) with the convex envelope based constraints (8), as detailed in Remark 3.4.
  • $\mathsf{ML}$: This is the mixed-integer linear formulation based on multilinear extension of attraction value functions, and its deviation is detailed in Appendix 11.1.
  • $\mathsf{SCIP}$: This refers to solving the binary nonlinear formulation (HAP) using the global optimization solver SCIP.
We compare the performance of the four formulations on synthetically generated problem instances. We generate these instances as follows. For a fixed number of products $N$ and memory length $M$, the base utility of each product $\beta_{i}^{0}$ is drawn from a uniform $[-1,1]$ distribution and its revenue $r_{i}$ from a uniform $[1,10]$ distribution. The no-purchasee utility is 00. For each $i\in[N]$ and $m\in[M]$, we set the history-dependent effect $\beta_{i}^{m}$ from two uniform distributions $U[-1,0]$ and $U[-2,-1]$ to represent a weak and strong satiation effect, denoted as $W$ and $S$, respectively. We vary $N\in\{30,50,70\}$, $M\in\{1,2\}$, and $T\in\{5,10\}$. For each combination of $N$, $M$, and a satiation level (W or S), we generate five instances of $\{r_{i},\beta_{i}^{0},\beta_{i}^{m}\}_{i\in[N],m\in[M]}$.
We solve each instance of (HAP) using one of the formulations $\mathsf{Conic}$, $\mathsf{Env}$, $\mathsf{ML}$, and $\mathsf{SCIP}$. We report the computation time and the end gap, which is defined as
$G_{\textsf{end}}=100\%\times\frac{R_{\mathsf{U}}-R_{\mathsf{IP}}}{R_{\mathsf{IP}}},$
where $R_{\mathsf{U}}$ and $R_{\mathsf{IP}}$ is the best upper bound and the best integer solution at termination with a time limit of 3600s and an optimality gap tolerance of $0.5\%$, respectively.
Table 4: Formulation efficiency comparison
Conic Env ML SCIP
T M N $\boldsymbol{\beta}$$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$
5 1 30 W 5 36 0 5 6 0 5 267 0 0 - 648
5 1 30 S 5 90 0 5 96 0 5 527 0 0 - 920
5 1 50 W 3 65 0.64 5 657 0 3 473 1.04 0 - 1380
5 1 50 S 3 1766 0.66 4 1249 0.68 1 3492 1.51 0 - 1527
5 1 70 W 3 453 1.75 4 102 2.17 3 855 2.26 0 - 1798
5 1 70 S 1 1695 0.87 2 757 0.68 1 2523 1.87 0 - 2191
5 2 30 W 5 267 0 5 95 0 3 1339 0.99 - - 913
5 2 30 S 3 527 2.39 3 363 1.5 0 - 11.76 0 - 1046
5 2 50 W 3 689 0.95 5 160 0 1 2760 1.57 0 - 1415
5 2 50 S 0 - 2.35 1 2133 1.64 0 - 49.13 0 - 1756
5 2 70 W 0 - 1.81 3 811 1.6 0 - 16.35 0 - 2369
5 2 70 S 1 1020 3.41 1 57 2.37 0 - 79.22 0 - 2459
10 1 30 W 0 - 0.92 4 422 0.86 0 - 0.94 0 - 744
10 1 30 S 0 - 2.14 1 2712 1.77 0 - 2.86 0 - 1005
10 1 50 W 0 - 1.35 2 1994 1.4 0 - 1.48 0 - 1477
10 1 50 S 0 - 3.82 0 - 3.12 0 - 3.88 0 - 1577
10 1 70 W 0 - 2.44 1 313 2.06 0 - 3.27 0 - 1907
10 1 70 S 0 - 3.08 0 - 1.9 0 - 4.43 0 - 2305
10 2 30 W 0 - 2.37 0 - 1.4 0 - 4.15 0 - 1001
10 2 30 S 0 - 7.52 0 - 5.8 0 - 165.07 0 - 1151
10 2 50 W 0 - 2.9 0 - 2 0 - 63.85 0 - 1563
10 2 50 S 0 - 8.73 0 - 6.05 0 - 608.16 0 - 1891
10 2 70 W 0 - 4.66 0 - 3.05 0 - 283.46 0 - 2756
10 2 70 S 0 - 6.51 0 - 4.64 0 - 750.15 0 - 2763
[…]
Besides the computation time and end gap, it is also helpful to compare the continuous relaxation of different formulations. We compute the root gap as follows
$G_{\textsf{root}}=100\%\times\frac{R_{\textsf{Rlx}}-R_{\textsf{IP}}}{R_{\textsf{IP}}},$
where $R_{\mathsf{Rlx}}$ is the objective value of the continuous relaxation of a given formulation. A small root gap indicates a tight formulation. Table 5 reports the average root gap of Conic, Env, and ML under different parameter configurations. Overall, the root gap of our formulations Conic and Env is significantly smaller than that of ML. It indicates the tightness of our formulations. Moreover, the root gap of our formulations is stable and does not scale with the problem size. For instance, the root gap of Env is no more than $6\%$ when $M=1$ and no more than $16\%$ when $M=2$.
[…]
T M N $\boldsymbol{\beta}$ Conic Env ML SCIP T M N $\boldsymbol{\beta}$ Conic Env ML SCIP
5 1 30 1 3.86 3.22 156.73 - 10 1 30 1 4.38 3.65 179.23 -
5 1 30 2 6.81 5.89 301.08 - 10 1 30 2 7.83 6.56 343.77 -
5 1 50 1 2.89 2.5 295.99 - 10 1 50 1 3.3 2.82 336.43 -
5 1 50 2 5.64 5.04 517.19 - 10 1 50 2 7.11 5.64 588.26 -
5 1 70 1 3.95 3.55 400.53 - 10 1 70 1 4.61 4.77 455.25 -
5 1 70 2 3.84 3.45 713.79 - 10 1 70 2 5.05 4.2 810.67 -
5 2 30 1 4.57 4.58 224.85 - 10 2 30 1 5.64 5.61 275.29 -
5 2 30 2 13.17 13.17 587.57 - 10 2 30 2 15.4 15.32 786.57 -
5 2 50 1 4.77 4.77 430.75 - 10 2 50 1 5.9 5.76 539.85 -
5 2 50 2 10.1 10.06 921.6 - 10 2 50 2 13.34 12.09 1419.53 -
5 2 70 1 5.47 5.3 717.98 - 10 2 70 1 6.94 6.46 940.42 -
5 2 70 2 6.85 6.43 1243.68 - 10 2 70 2 9.02 7.85 1890.09 -
6.2 Computing Two-Cyclic Policies
In this section, we demonstrate the efficiency and scalability of (11) for computing two-cyclic policies. We propose a projected cutting plane algorithm to implement (11). Note that due to the presence of $\boldsymbol{\Gamma}$ variables, the number of variables in (11) is $O(N^{2})$. To address this issue, we propose a projected cutting plane algorithm to generate a compact approximation of (11). The cutting plane algorithm first sets the linearized (2-Cyclic) as a base formulation and obtains a fractional solution using the continuous relaxation of the base formulation. The cutting plane algorithm then uses a separation oracle to generate new cuts, which are the projection of constraints (11c)-(11e) violated by the fractional solution to the space of $(\boldsymbol{x},\boldsymbol{\gamma})$ variables. Adding new cuts into the base model leads to a formulation with a tighter continuous relaxation. Running the separation oracle once is a round of cut generation. We can add new cuts for $K$ rounds and the final formulation as BF-$K$, where BF denotes “bound-free”. Indeed, one can repeat this process (i.e., a large $K$) until the fractional solution of BF-$K$ does not violate (11c)-(11e). We observe that the marginal improvement of formulation tightness is decreasing. Hence, we can set a small $K$ and obtain a much tighter formulation than the base model. More specifically, we set $K=1$ to compute the two-cyclic policy because BF-$1$ has the most marginal improvement of the root gap. Recall that a small root gap means a tight continuous relaxation. The relative improvement in the root gap of BF-$K$ compared to the conic formulation (Cycle-Conic) with non-overlapping constraint is
$G_{\textsf{root}}^{K}=100\%\times\frac{G_{\textsf{root,conic}}-G_{\textsf{root,BF-K}}}{G_{\textsf{root,conic}}},$
where $G_{\textsf{root,BF-K}}$ and $G_{\textsf{root,conic}}$ are the root gap of BF-$K$ and (Cycle-Conic), respectively.
Table 6: The computational performance of (11) with a cutting plane implementation
N $\boldsymbol{\beta}$ $G^{1}_{\mathsf{root}}(\%)$ $G^{2}_{\mathsf{root}}(\%)$ $G^{3}_{\mathsf{root}}(\%)$ $T_{\mathsf{BF}-1}$(s)
500 W 98.36 98.54 98.66 426
500 S 98.53 98.71 98.80 523
750 W 98.67 98.83 98.90 579
750 S 98.57 98.74 98.83 604
1000 W 98.79 98.89 98.96 606
1000 S 98.90 99.01 99.09 611
Table 6 summarizes the performance of different formulations. BF-$1$ has the most marginal improvement of the root gap. After adding one-round cuts, the natural continuous relaxation of BF-$1$ approximately shrinks $98\%$ root gap of the conic formulation (Cycle-Conic). It suggests that the natural continuous relaxation of BF-$1$ is extremely close to the optimal integer solution and much tighter than the continuous relaxation of (Cycle-Conic), which supports our theoretical result in Theorem 4.4. The last column reports the average running time, including the time for cut-generation, of using formulation BF-$1$ to compute an optimal two-cyclic policy. The average computation time for large instances (e.g., $N=1000$) is about $600$ seconds. This shows the computational efficiency and scalability of the bound-free formulation (11).
7 Conclusion
This paper studies a multi-period assortment planning problem in which customer utility depends on historical assortments. We first show that the problem is solvable in polynomial time if and only if customers’ utility increases with historical repeated offerings. The optimal policy is a sequential revenue-ordered policy. However, the problem becomes NP-hard if the historical effect is the opposite, even with a planning horizon of two periods. We develop an MIECP formulation to find a global optimal solution. Our numerical analyses indicate that our formulation efficiently achieves optimality and generates higher revenue than heuristic policies, especially under strong satiation effects. We further explore cyclic policies when the problem has a large time horizon. We theoretically prove that a cyclic policy with a cycle equal to the memory length plus one is asymptotically optimal. Additionally, we develop a neat, bound-free reformulation for non-overlapping cyclic policies when the memory length equals one.
Our research can be extended in several directions. The first direction would be to explore the joint assortment and pricing or inventory problem under history-dependent utility. The second direction would be to determine, for online settings under the history-dependent effect, how to provide personalized assortments when customers stochastically revisit the online store. Third, it would be interesting to examine how to incorporate history-dependent effects into different choice models and optimize assortment decisions.