### v3

# Assortment Optimization Under History-Dependent Effects

[…]
Assortment Optimization Under History-Dependent Effects
[…]
This paper examines how to plan multi-period assortments when customer utility depends on historical assortments. We formulate this problem as a nonlinear integer programming model and show it is NP-hard in the presence of a negative history-dependent effect (such as a satiation effect). We build solution methodologies for obtaining global optimal solutions under a general setting that the history-dependent effects could be a mixture of positive and negative. We propose using a lifting-based framework to reformulate the problem as a mixed-integer exponential cone program that state-of-the-art solvers can solve. We also design a sequential revenue-ordered policy and show that it solves our problem to optimality in polynomial time when historical assortments positively affect customer utility (such as an addiction effect). Additionally, we identify an optimal cyclic policy for an asymptotic regime, and we also relate its length to the customer’s memory length. Finally, we present a case study using a catering service dataset, showing that our model demonstrates good fitness and can effectively balance variety and revenue.
[…]
1 Introduction
Many companies sell products to customers who make repeat purchases or visit over multiple periods. For example, corporate cafeterias provide workday luncheons for employees; online grocery flash sell platform plans assortments for customers who repeatedly visit the online store and seek deep discounts. Customers’ utility may vary over time and may divert to products offered by competitors if they feel “bored”. Therefore, these companies must carefully plan their daily assortments over time to keep their customers satisfied.
A typical example is the corporate dining industry. Corporate dining is an essential but often overlooked industry. It was reported that a working person will spend an average of \$2500 per year on working-day lunches (ezCater 2024). Though corporate cafeterias are convenient and usually offer subsidized prices, they often face complaints of boredom due to the need for more dish variety. To address this, cafeteria managers manually diversify their menus based on their experience and industry knowledge. However, without a careful understanding of how previous menus impact customer demand, manual changes may be inefficient.
Many studies have explored and validated that customers’ historical experiences may affect their willingness to buy or retain a product (Bowden 2009, Lemon and Verhoef 2016). For example, a customer may initially feel satisfied with a product, but this satisfaction may wane over time. This utility decay phenomenon is also observed in fashion products (Caro et al. 2014). On the other hand, a customer may also prefer to consistently consume “traditional and familiar” food that could maintain or increase her utility (Edenred 2024). Today, with the increasing collection of transactional data, companies seek to leverage these data to estimate customers’ utility changes and build data-driven assortment planning systems. For instance, new technology companies have emerged that revolutionize traditional catering services by analyzing transactional data to estimate customers’ preferences and offering optimized daily cafeteria menus (Fooda 2024).
Besides corporate cafeterias, analogous situations arise in various other sectors. Flash sales platforms must organize their promotional assortments ahead of time, based on the products suppliers provide (de Albéniz et al. 2020). For instance, Meituan, a leading Chinese online shopping platform, manages a grocery flash sale channel. Due to the lead time of contract negotiations with suppliers, Meituan needs to decide, in advance, assortments for daily deals over a fixed planning horizon. Given that customers frequently visit the channel for discounted groceries, planning must consider historical offerings. More details about Meituan’s flash sale campaign are in the ecompanion EC.1. Managers of cafeterias and flash sales encounter a common multi-period planning problem that decides how to offer identical assortments to all customers over a fixed planning horizon. In practice, these assortments are usually predetermined before the planning horizon, possibly due to limits (such as lead time) from the supplier side.
In this paper, we focus on the above applications and investigate the corresponding multi-period assortment planning problem considering history-dependent utilities to maximize the long-run average revenue. The history-dependent effects bring new challenges for assortment planning because of the interplay between cross-product and cross-period effects. Cross-product effects refer to the influence of other products in an assortment on purchase decisions. For instance, customer substitution behavior is commonly modeled using discrete choice models, like the multinomial logit (MNL) model (McFadden 1974). Cross-period effects refer to the temporal influences on customer utility, i.e., the utility of a product in a given period depending on its offering history. Given both types of effects, the offering decisions for multiple products over multiple periods are made simultaneously. The number of candidate assortments, therefore, scales exponentially with the number of products and the planning horizon. This can be enormous for even a practically-sized problem. Moreover, customer behavior under both types of effect makes the decision more complex because, for a given period, the purchase probability is a nonlinear function—specifically, a softmax function—of the history-dependent utility functions.
[…]
The average out-of-sample results of three assortment plans Our model is the history-dependent model. We normalize the revenue generated by the current practice to one. The assortment size is the number of product types offered in the planning horizon.
Despite the complexity of assortment planning with history-dependent effects, the benefit of considering such effects is significant. Using the data from a corporate cafeteria with which we cooperated, we evaluate assortments derived from a history-dependent model (our model, estimated in Section 5), from the MNL model, and the current practice observed in the data (as unit one for normalization). The results shown in Figure 1 indicate that our model achieves the highest revenue in the out-of-sample test, about $10.4\%$ more than the revenue generated by the MNL model. Our model also produces assortments that have greater variety than those obtained by the MNL model. The current practice decided by managers resulted in assortments with a greater variety but the lowest revenue. See Section 5 for details of the case study.
1.1 Description of Results
We study a multi-period assortment planning problem in which a firm facing both cross-product and cross-period effects needs to decide what to offer in each period to maximize the average expected revenue. We model customer choice in a given period using a variant of the MNL model in which the utility of each product linearly depends on its past offering decisions. In our model, an assortment decision for a given period affects both current and future purchase behavior in a nonlinear fashion. In other words, our assortment planning problem is a nonlinear integer programming problem.
We show that this problem is NP-hard under the presence of negative history-dependent effects and propose exact solution methodologies for a general setting that the history-dependent effects could be a mixture of addiction and satiation. We propose a lifting-based reformulation framework to obtain a mixed-integer exponential cone program (MIECP) formulation. The key idea behind our formulation is to lift the original nonconvex constraints into a higher-dimensional space where the lifted constraints are simpler to model. This reformulation allows us to utilize prevalent techniques in the optimization literature, such as the McCormick envelope, the perspective formulation, and the Lovász extension, to model the lifted structures. Numerical studies indicate that our formulation efficiently achieves optimality and yields higher revenue than heuristic policies. We also identify an optimal sequential revenue-ordered (RO) policy if history-effects are all positive, i.e., addiction effects. This characterization indicates that our problem with the addiction effect is solvable in polynomial time. For problems with a long-term planning horizon, theoretically, we show that a cyclic policy is optimal and find that, under a certain non-overlapping condition, the optimal cycle length equals the customer memory length plus one. Computationally, finding such policies is NP-hard but also admits an MIECP formulation. Moreover, the non-overlapping condition leads to a mixed-integer linear programming (MILP) formulation that is stronger than the conic one.
We conduct a case study using transactional data from a corporation cafeteria operated by a high-tech catering service company in China. The cafeteria primarily serves workday lunches for employees who work nearby. We estimate customers’ utility function and then design weekly day-by-day lunch menus. The estimation shows that a satiation effect does exist, but it only lasts for a short period (i.e., two or three days). We demonstrate that our model can provide menus that balance variety and revenue, while current industry practices sacrifice revenue, and the classic MNL model yields menus with limited variety and low revenue.
We highlight the contributions of our paper as follows:
  • 1.
    Our model, which we refer to as HAP (i.e., history-dependent assortment problem), simultaneously incorporates the cross-product and cross-period effects into the multi-period assortment planning process. The cross-period history-dependent effects could be all positive, all negative, and mixed across products and periods. We demonstrate, via a case study of a corporate cafeteria, that our model has several benefits. First, our result suggests that the service provider should provide a greater variety of products if satiation effects exist. When satiation effects last longer, it is beneficial to offer more types of products across periods. Furthermore, it is recommended that a product be offered only once within the memory length to prevent a utility decrease due to satiation effects, particularly in the case of products with high prices. This observation naturally leads to a “non-overlapping” cyclic assortment planning policy, which we explore in Section 4, where we establish its optimality. Most importantly, our history-dependent policy effectively balances variety and the revenue objective, resulting in assortments that achieve the highest revenue in out-of-sample tests while maintaining the same variety as found in practical scenarios.
  • 2.
    We characterize, in terms of history-dependent effects, the conditions under which a sequential revenue-ordered policy globally solves our HAP model. In general, it can be shown that finding a global optimal solution is NP-hard. To tackle this difficulty, we develop an MIECP formulation of our HAP model by using lifting and coupling ideas from the optimization literature, including perspective formulation, the McCormick envelope, and the Lovász extension. In particular, we modify the Lovász extension to model attraction values under general history-dependent effects. Our formulations, together with cutting-plane algorithms, outperform an MILP formulation of HAP obtained using a prevalent strategy in the mixed-integer nonlinear programming (MINLP) literature and the original nonlinear integer formulation HAP solved by the state-of-the-art MINLP solver SCIP.
  • 3.
    We establish the asymptotic optimality of cyclic policies, and finding such policies allows an MIECP formulation. Our study reveals a connection between the length of an optimal cyclic policy and the length of customer memory. In particular, we characterize the condition under which the optimal cyclic length is the customer memory length plus one. This condition, in addition, leads to an MILP formulation for finding optimal cyclic policies, which is, theoretically, tighter than the previous conic formulation.
1.2 Literature Review
Extensive research has incorporated history-dependent effects into operations problems, including pricing and promotion (Bi et al. 2025, Chen et al. 2023c), service design (Bernstein et al. 2022, Lei et al. 2023), resource allocation (Adelman and Mersereau 2013), retention management (Kanoria et al. 2024), etc. Studies show that ignoring history-dependent effects may lead to substantial losses. For instance, considering the reference price can reduce $5.82\%$ prediction error of sales (Wang 2018), and the optimal bundle accounting for customer satiation improves the revenue by more than $4.5\%$ (Gürlek et al. 2023).
History-dependent effects are receiving increasing attention in assortment planning. One stream of work is to design personalized assortment planning for online platforms. The history-dependent effect derives from repeated interactions between the firm and customers, which breaks the traditional single-interaction assumption in personalized assortment optimization literature (El Housni and Topaloglu 2023). Based on the easily accessible purchase history of a customer, these studies consider two settings: (i) customers have stable repeated interactions with the firm, e.g., Chen et al. (2023b) optimize adaptive assortment decisions based on information inferred from a customer’s historical purchases; (ii) a customer can leave or churn if he or she is not satisfied. Hence, many studies focus on designing delicate exploration-exploitation recommendation algorithms that may limit exploration budgets (Bastani et al. 2022, Sumida and Zhou 2023).
For non-personalized assortment planning, the history-dependent effect is captured by a dynamic choice/demand function of the entire market. The effect may derive from attributes of a product itself, e.g., declining utility. Such intertemporal connections lead to a nonlinear and high-dimensional problem. Two settings are considered: (i) an adaptive assortment policy, e.g., Caro and Martínez-de Albéniz (2020) consider a content release problem where the attraction of content decays and the number of followers stochastically evolves; (ii) a non-adaptive assortment plan for the entire horizon made before the first period, e.g., Caro et al. (2014) optimize release schedule of fast fashion products with exponentially declining attraction before period one due to a long lead time. Our work contributes to the second type of literature. Our model can represent the model in Caro et al. (2014) and adds more flexibility (see details in Section 2.2), which brings new challenges and needs additional methods.
Our paper contributes to the growing literature on developing new solution methods for assortment planning. The three primary approaches include the revenue-ordered policy, approximation algorithms, and integer programming for obtaining exact solutions. Since Talluri and Van Ryzin (2004) showed that the optimal assortment under the MNL model is revenue-ordered, this policy has been extensively studied in various contexts (Wang and Wang 2017, Sumida and Zhou 2023). For example, Gao et al. (2021) prove that the revenue-ordered assortment policy is optimal under the MNL model when the offered assortment is gradually revealed. Xu and Wang (2023) exploit the revenue-ordered property to solve a multistage assortment optimization problem. We contribute to this literature by showing that a sequential revenue-ordered policy is optimal for our problem under the addiction effect.
A stream of studies focuses on designing approximation algorithms for assortment plannings. For instance, Feldman and Topaloglu (2015) and Désir et al. (2020) design constant factor approximation algorithms for capacitated assortment optimization under the nested logit model and the Markov chain choice model, respectively. Later, Désir et al. (2022) propose a fully polynomial time approximation scheme (FPTAS) for capacitated assortment problems under the mixed MNL, Markov chain, and nested logit choice models. FPTAS are also proposed to solve assortment optimization under the exponential choice model (Aouad et al. 2023), with online dynamic behavior (Liu et al. 2020, Feldman and Jiang 2023, Aouad et al. 2024), and with multi-purchase behavior (Jasin et al. 2024, Luan et al. 2025).
[…]
Last, we comment that our paper contributes to a growing set of applications of conic programming/discrete optimization in designing operations systems, including joint inventory-location problems (Atamtürk et al. 2012), battery swap networks design (Mak et al. 2013), service delivery scheduling (Kong et al. 2013), process flexibility design (Simchi-Levi and Wei 2012, Yan et al. 2018), electric vehicle charging (Chen et al. 2023a), proactive policing (He et al. 2023), network flow (Simchi-Levi et al. 2019), and nonlinear resource allocation (He and Tawarmalani 2024).
1.3 Structure
The rest of the paper is organized as follows. In Section 2, we formally define the assortment planning problem and establish its computational complexity. In Section 3, we propose a mixed-integer exponential cone formulation that can handle any mixed addiction-satiation history-dependent effects. We prove that, under positive effects, a revenue-ordered policy is optimal. In Section 4, we present our results on cyclic policies. We present a case study using real data in Section 5 and report numerical results on synthetic data in Section 6. We conclude the paper in Section 7. All proofs are given in the ecompanion.
2 Model
In Section 2.1, we define our multi-period assortment planning problem with history-dependent effects and establish its computational complexity. In Section 2.2, we discuss the model features in detail.
2.1 Incoporating History-Dependent Effects into Choice Models
Consider a firm that sells $N$ products to customers. Let $[N]:=\{1,\ldots,N\}$ denote the set of all products, and let $[N]^{+}:=[N]\cup\{0\}$, where the index 00 denotes the outside option. For each product $i\in[N]$, let $r_{i}$ denote the revenue obtained from selling an instance of product $i$. There is a total of $T$ planning horizons, and the firm needs to decide what to offer in each period. More specifically, for each period $t\in[T]$, let $\boldsymbol{x}^{t}=(x^{t}_{1},\ldots,x^{t}_{N})\in\{0,1\}^{N}$ be a binary decision variable modeling the subset of products offered in period $t$, i.e., $x^{t}_{i}=1$ if and only if product $i$ is offered in period $t$. We assume that in each period, one unit mass of customers arrives, and the customers select an item from the offered products or leave without a purchase. For example, in a setting such as corporate cafeterias, the firm serves relatively stable customers (e.g., most customers are employees who work nearby), and customers make purchase decisions on a daily basis.
We modify the MNL model to encapsulate the impact of the historical assortments on customer choices. Each product has a utility $U_{i}$, which depends on attributes of the product and environment. Following Guadagni and Little (1983), Wang (2018) and Xie and Wang (2024), we assume that the utility of one product is a linear function of its attributes and historical offerings. Specifically, the utility of product $i$ in period $t$ affinely depends on its historical offering records within a memory length $M$:
$U_{i}^{t}=\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta_{i}^{m},$ (Utility)
where $M\in\operatorname{\mathbb{Z}}_{+}$ denotes the customer memory length. For simplicity in the presentation, we assume throughout this paper that $x_{i}^{k}:=0$ for $k\leq 0$. The parameter $\beta^{0}_{i}$ is the base utility of the product $i$ determined by its attributes. The parameter $\beta_{i}^{m}$ is the incremental utility if product $i$ is offered in the $m^{\text{th}}$ period before the current time. Let $\boldsymbol{\beta}:=(\beta_{i}^{m})_{i\in[N],m\in[M]}$ denote the history-dependent effect matrix. The value of $\beta_{i}^{m}$ can be positive or negative. A positive $\beta_{i}^{m}$ indicates that exposure to product $i$ increases its utility, acting as a stimulus (Fox et al. 1997). While a negative $\beta_{i}^{m}$ suggests a satiation effect, leading to customer fatigue from repeated offerings. The history-dependent effects can vary across products and periods. In our planning model, the assortments are fixed for each period. Therefore, we do not consider personalized assortments or model individual-level utility functions.
For a given assortment plan $\boldsymbol{x}:=(\boldsymbol{x}^{1},\cdots,\boldsymbol{x}^{T})$, we use $\pi^{t}_{i}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ to denote the purchase probability of product $i$ in period $t$. Under the MNL model, the purchase probability of product $i$ in period $t$ is
$\pi^{t}_{i}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ $=\dfrac{x^{t}_{i}\exp\bigl{(}\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta^{m}_{i}\bigr{)}}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}}\quad\operatorname{for}i\in[N],$
$\pi^{t}_{0}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ $=\dfrac{1}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}},$
where we normalize the utility of the outside option to 0. The purchase probabilities depend on two components: the current assortment $\boldsymbol{x}^{t}$ and the past assortments $\boldsymbol{x}^{t-1},\dots,\boldsymbol{x}^{t-M}$. The former dependence captures cross-product effects such as the substitution effect, while the latter dependence models cross-period effects. The purchase probability can also be viewed as the market share of product $i$ in period $t$ and presents the effect of historical assortments on sales of the entire market (Batsell and Polking 1985).
With these definitions, assortment planning under history-dependent effects can be expressed as the following fractional binary program, called the history-dependent assortment problem:
$\max\Biggl{\{}\sum_{t\in[T]}\frac{1}{T}\sum_{i\in[N]}r_{i}\pi^{t}_{i}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})\Biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\,\operatorname{for}t\in[T]\text{ and }\boldsymbol{x}\in\mathcal{P}\Biggr{\}},$ (HAP)
where $\mathcal{X}$ (resp. $\mathcal{P}$) is a system of linear inequalities that models cross-product constraints (resp. cross-period constraints). The cross-product constraints often include resource restrictions within a single period, such as cardinality and capacity constraints, and address various operational considerations in the assortment planning literature (Sen et al. 2018, Chen et al. 2024). Cross-period constraints describe the intertemporal relationships of assortments across different periods, such as non-overlapping assortments over consecutive periods, which occur in multi-period assortment planning problems (Liu et al. 2020, Chen et al. 2023b).
This model is generally difficult to solve. We show in the following proposition that  (HAP) is NP-hard under very simple settings if history-dependent effects are negative.
Proposition 2.1
(HAP) is NP-hard even when the planning horizon is two, the memory length is one, and history-dependent effects are negative.
2.2 Model Discussion
We outline here the main features of our model, justify our assumptions, and identify relevant scenarios.
It is apparent that our formulation requires a firm to present the same assortment to all customers in each period before the planning horizon begins. In other words, the firm is unable to take advantage of personalized assortments and adaptive policies, which are common practices in online assortment optimization such as Amazon’s recommendation system. However, those practices are not suitable for the case of corporate cafeterias, where it is common to treat employees with the same menu, and it is impractical to dynamically adjust assortments due to the lead time in both procurement and cooking processes (Van Wezel et al. 2006). Observations from our partnered catering company indicate that the corporate cafeteria plans the lunch menus on a weekly basis. Moreover, our model assumption is also reasonable for flash sale channels (i.e., predetermine assortments for several days due to the negotiation time with suppliers).
Our model describes the history-dependent effect through a linear utility function with historical offerings as features. The linear utility function assumption is common in the logit-based choice model to capture impacts such as product attributes (Guadagni and Little 1983) and historical prices (Wang 2018). Moreover, the linear utility function covers multiplication in the attraction value in literature. For example, our model with negative history-dependent effects represents the model in Caro et al. (2014). Our model offers additional flexibility by allowing for positive, negative, and mixed history-dependent effects and offering products in any period.
In the linear utility function, we incorporate the historical offerings to capture the history-dependent effect. The utility function captures the direct effect of historical assortments that are critical decisions at a firm level. Moreover, we only consider the direct history-dependent effect, i.e., the historical offering of a product will only affect the product’s future utility itself. We do not include the cross-product history-dependent effect, i.e., the historical offering of a product might affect the utility of other products, because this effect is usually at a much smaller scale (e.g., Vilcassim et al. 1999, Dubé and Manchanda 2005, Mehta 2007). In our empirical estimation, the cross-product history-dependent effect is approximately one-tenth of the direct effect and is significant only when $M=1$. We believe that our model efficiently captures the primary effect and maintains model tractability.
We manage the history-dependent effects across periods through the memory length $M$, inspired by the patience level from dynamic pricing studies (Popescu and Wu 2007, Liu and Cooper 2015). This patience level indicates how many periods a customer will wait for a price reduction. Consequently, it is logical to assume that customers might retain a limited memory length for a product.
Our model does not include inventory decisions. In cafeteria settings, for instance, procurement choices can be made subsequent to assortment planning. Additionally, inventory rollover is not considered in the study because cooked dishes cannot be kept for the following day due to hygiene standards. To simplify the analysis, this study will concentrate on the assortment planning problem with cross-product and cross-period effects, which is already a challenging problem.
3 Solution Approaches
In this section, we aim to solve the history-dependent assortment problem (HAP) to its global optimality. In Section 3.1, we propose a mixed-integer exponential cone formulation for the general case, and, in Section 3.2, we characterize conditions under which a sequential revenue-ordered policy is optimal.
3.1 Integer Programming Approach
Due to the combinatorial and nonconvex nature of (HAP), state-of-the-art global optimization solvers such as BARON (Tawarmalani and Sahinidis 2005) and SCIP (Bestuzheva et al. 2021) fail to solve practical-sized instances—see Section 6.1 for computational evidence. Nevertheless, we focus on developing techniques for obtaining mixed-integer programming formulations of (HAP), which will allow us to leverage modern mixed-integer program solvers such as GUROBI and MOSEK. Then, in Section 4, we will use these techniques to find asymptotically optimal cyclic policies.
Our formulation introduces the following additional variables to represent the customer choice behavior. For each period $t\in[T]$, we use $\rho^{t}$ to denote the no-purchase probability for period $t$:
$\rho^{t}=\frac{1}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}},$
and for each product $i\in[N]$, we use $y^{t}_{i}$ to denote the choice probability for product $i$ in period $t$:
$y^{t}_{i}=\frac{x^{t}_{i}\exp\bigl{(}\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta^{m}_{i}\bigr{)}}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}}.$
In what follows, we refer to $\rho^{t}$ as the no-purchase probability variable and $y^{t}_{i}$ as the choice probability variable. Using these definitions, we can reformulate  (HAP) as follows:
$\max_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{\rho}}\quad$ $\frac{1}{T}\sum_{i\in[N]}\sum_{t\in[T]}r_{i}y_{i}^{t}$
$\operatorname{s.t.}\quad$ $\boldsymbol{x}^{t}\in\{0,1\}^{N}\cap\mathcal{X}\text{ and }\boldsymbol{x}\in\mathcal{P}$ $\operatorname{for}\,t\in[T]$
$\rho^{t}+\sum_{i\in[N]}y_{i}^{t}=1\quad$ $\operatorname{for}\,t\in[T]$
$y^{t}_{i}=\rho^{t}x^{t}_{i}\exp\Bigl{(}\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta^{m}_{i}\Bigr{)}\quad$ $\operatorname{for}\,t\in[T]\text{ and }i\in[N].$ (Choice)
The objective function is a non-negative weighted combination of choice probability variables. The second constraint ensures that the sum of the choice probability and the no-purchase probability equals one for each period. The last constraint represents the choice probability for product $i$ in period $t$, and thus we refer to it as constraint  (Choice). This reformulation encapsulates the nonconvexity of problem  (HAP) in constraint (Choice). Hence, it suffices to derive a convex representation of constraint (Choice).
The main component of constraint (Choice) is the composition of a univariate exponential function and an affine utility function, defined as follows:
$\alpha_{i}(z^{1},\ldots,z^{M}):=\exp\biggl{(}\beta_{i}^{0}+\sum_{m\in[M]}\beta^{m}_{i}\cdot z^{m}\biggr{)}\qquad\operatorname{for}\;(z^{1},\ldots,z^{M})\in\{0,1\}^{M}.$
For a given assortment plan $\boldsymbol{x}:=(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{T})$, we interpret $\alpha_{i}(x_{i}^{t-1},\ldots,x_{i}^{t-M})$ as the attractiveness of product $i$ in period $t$. Therefore, we will henceforth refer to $\alpha_{i}(\cdot)$ as the attraction value function. To handle the discrete nature of the attraction value function, we will use the notion of a continuous extension. Given a function $f:\{0,1\}^{n}\to\mathbbm{R}$, a function $g:[0,1]^{n}\to\mathbbm{R}$ is called a continuous extension of $f(\cdot)$ if $g(\cdot)$ is continuous and $g(\boldsymbol{z})=f(\boldsymbol{z})$ for every $\boldsymbol{z}\in\{0,1\}^{n}$. In addition, if an extension function $g(\cdot)$ is convex (resp. concave), we say that it is a convex (resp. concave) extension. It is common for a discrete function to admit many convex (resp. concave) extensions. A natural idea is to use the tightest convex/concave extension. Since the discrete domain $\{0,1\}^{n}$ is the vertex set of the continuous domain $[0,1]^{n}$, the pointwise largest convex (resp. smallest concave) extension of $f(\cdot)$ is the convex (resp. concave) envelope, denoted as $\operatorname{conv}(f)(\cdot)$ (resp. $\operatorname{conc}(f)(\cdot)$), of $f(\cdot)$ over the box $[0,1]^{n}$ (see Tawarmalani and Sahinidis 2002, Theorem 6).
3.1.1 A Lifted Representation of (Choice)
In this subsection, we focus on modeling constraint (Choice). The idea is to lift the original nonconvex constraint (Choice) into a higher-dimensional space where the lifted constraint is dramatically simpler to represent than the original constraint. Specifically, this will allow us to invoke prevalent techniques, such as the McCormick envelope (McCormick 1976), the perspective formulation (Günlük and Linderoth 2010), and the Lovász extension (Lovász 1983), to model the lifted nonconvex structures.
[…]
$\lambda+\lambda f\Bigl{(}\frac{\lambda\cdot\boldsymbol{x}}{\lambda}\Bigr{)}\leq 1,$
where $\lambda f(\frac{\lambda\cdot\boldsymbol{x}}{\lambda})=0$ when $\lambda=0$. Next, we replace $\lambda\boldsymbol{x}$ with a new variable vector $\boldsymbol{u}$. This leads to
$\boldsymbol{u}=\lambda\boldsymbol{x}\quad\text{and}\quad\lambda+\lambda f\Bigl{(}\frac{\boldsymbol{u}}{\lambda}\Bigr{)}\leq 1,$
where the first constraint can be linearized using the McCormick envelope McCormick (1976) and the second constraint is convex, since the function $\lambda f(\boldsymbol{u}/\lambda)$ is the perspective function of the convex function $f(\cdot)$ (Rockafellar 1970). Formally, given a function $f:D\to\mathbbm{R}$, where $D$ is a bounded subset of $\mathbbm{R}^{n}$, the perspective function of $f$, denoted as $\operatorname{pers}(f)$, is defined as
$\operatorname{pers}(f)(\lambda,\boldsymbol{x})=\begin{cases}\lambda f(\frac{\boldsymbol{x}}{\lambda})&\text{ for }\lambda>0\text{ and }\boldsymbol{x}\in\lambda\cdot D\\ 0&\text{ for }\lambda=0\text{ and }\boldsymbol{x}=\boldsymbol{0}.\end{cases}$
Perspective reformulations have been widely used in solving nonconvex optimization problems (Günlük and Linderoth 2010); see Bestuzheva et al. (2023) for a review of applications and a detailed computational study of perspective reformulations.
We now present our lift of the constraint (Choice). Note that the choice probability of product $i$ in period $t$ is computed by scaling its attraction value function $\alpha_{i}(\cdot)$ with a nonlinear factor $\rho^{t}x_{i}^{t}$, that is,
$y^{t}_{i}=(\rho^{t}x^{t}_{i})\cdot\alpha_{i}(x_{i}^{t-1},\ldots,x_{i}^{t-M}).$
First, we introduce a new variable $\gamma_{i}^{t}$ to represent the nonlinear factor $\rho^{t}x_{i}^{t}$. Second, we use the new variable $\gamma_{i}^{t}$ to scale the historical assortment decisions $x^{t-m}_{i}$ with $m\in[M]$, and we introduce the variable $z_{im}^{t}$ to represent the scaled variable $\gamma_{i}^{t}x^{t-m}_{i}$. Using the additional variables $\gamma_{i}^{t}$ and $\boldsymbol{z}_{i}^{t}=(z_{i1}^{t},\ldots,z_{iM}^{t})$ and the definition of a perspective function, we decompose constraint (Choice) into constraints
$\gamma_{i}^{t}=\rho^{t}x_{i}^{t}$ (2a)
$z_{im}^{t}=\gamma_{i}^{t}x^{t-m}_{i}\quad\operatorname{for}\,m\in[M]$ (2b)
$y^{t}_{i}=\operatorname{pers}(\alpha_{i})(\gamma^{t}_{i},\boldsymbol{z}_{i}^{t}).$ (2c)
Constraint (2a) lifts the nonlinear scaling factor $\rho^{t}x_{i}^{t}$ to $\gamma_{i}^{t}$; constraint (2b) lifts the scaled historical assortment decision $\gamma_{i}^{t}x_{i}^{t-m}$ to $z_{im}^{t}$; and constraint (2c) perspectifies the attraction value function in the space of the newly introduced variables $\gamma_{i}^{t}$ and $\boldsymbol{z}_{i}^{t}$.
To obtain a mixed-integer convex formulation of constraint  (Choice) in the lifted space, we individually relax each of the three constraints in (2). Specifically, we use the McCormick envelope (McCormick 1976) to relax constraint (2a). To apply the McCormick envelope to our setting, we derive a lower bound $\rho^{t}_{L}$ and upper bound $\rho^{t}_{U}$ on the no-purchase probability variable $\rho^{t}$, and impose the following constraints:
$\gamma_{i}^{t}$ $\leq\min\bigl{\{}\rho^{t}_{L}x_{i}^{t}+\rho^{t}-\rho^{t}_{L},\ \rho^{t}_{U}x_{i}^{t}\bigr{\}}$ (3)
$\gamma_{i}^{t}$ $\geq\max\bigl{\{}\rho^{t}_{L}x_{i}^{t},\ \rho^{t}_{U}x_{i}^{t}+\rho^{t}-\rho^{t}_{U}\bigr{\}}.$
Similarly, using 00 (resp. $\rho^{t}_{U}$) as a lower (resp. upper) bound on $\gamma^{t}_{i}$, the McCormick envelope reformulates constraint (2b) as follows:
$z_{im}^{t}$ $\leq\min\bigl{\{}\gamma_{i}^{t},\rho^{t}_{U}x^{t-m}_{i}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M]$ (4)
$z_{im}^{t}$ $\geq\max\bigl{\{}0,\ \rho^{t}_{U}x^{t-m}_{i}+\gamma^{t}_{i}-\rho^{t}_{U}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M].$
Finally, we represent constraint (2c) by using two continuous extensions of the attraction value function $\alpha_{i}(\cdot)$. On one side, a natural continuous extension $\tilde{\alpha}_{i}:[0,1]^{M}\to\mathbbm{R}$ can be obtained from $\alpha_{i}(\cdot)$ by relaxing the binary assortment variables to continuous ones. This continuous function is convex since it is the composition of a convex function and an affine function. Using this extension, we relax constraint (2c) to
$y^{t}_{i}\geq\operatorname{pers}(\tilde{\alpha}_{i})(\gamma^{t}_{i},\boldsymbol{z}^{t}_{i}).$
This is a convex inequality since $\tilde{\alpha}_{i}(\cdot)$ is a convex function and so is its perspective function (Rockafellar 1970). Moreover, this constraint can be represented as an exponential cone constraint (Ben-Tal and Nemirovski 2001) (see more details in the ecompanion EC.3.2):
$y_{i}^{t}\geq\exp(\beta^{0}_{i})\cdot\gamma^{t}_{i}\cdot\exp\biggl{(}\frac{\sum_{m\in[M]}\beta_{i}^{m}z_{im}^{t}}{\gamma^{t}_{i}}\biggr{)},$ (5)
where the right-hand-side function can be regarded as a convex under-estimator of the choice probability variable $y^{t}_{i}$. On the other side, an exact representation of (2c) requires us to overestimate the choice probability variable. To do this, we replace $\alpha_{i}(\cdot)$ with its tightest concave extension or, equivalently, concave envelope. Using the concave envelope, we arrive at another convex relaxation of (2c):
$y_{i}^{t}\leq\operatorname{pers}\bigl{(}\operatorname{conc}(\alpha_{i})\bigr{)}(\gamma^{t}_{i},\boldsymbol{z}^{t}_{i}).$ (6)
Next, we show that an explicit linear description of this constraint can be obtained from perspectifying a variant of the Lovász extension of $\alpha_{i}(\cdot)$ (Lovász 1983).
The Concave Envelope of $\alpha_{i}(\cdot)$ Our construction is based on the Lovász extension of a set function. The Lovász extension is an extension of a function $f$ defined on $\{0,1\}^{n}$ to a function $\hat{f}$ defined on $[0,1]^{n}$. For each subset $S$ of $[n]$, let $\chi^{S}\in\{0,1\}^{n}$ be its indicator vector—that is, the $i^{\text{th}}$ coordinate of $\chi^{S}$ is $1$ if and only if $i\in S$. Observe that every vector $\boldsymbol{x}\in[0,1]^{n}$ can be expressed uniquely as $\boldsymbol{x}=\lambda_{0}\chi^{T_{0}}+\lambda_{1}\chi^{T_{1}}+\cdots+\lambda_{n}\chi^{T_{n}}$, where $\lambda_{k}\geq 0$ for $k=0,1,\ldots,n$ and $\sum_{k=0}^{n}\lambda_{k}=1$, and $\emptyset=T_{0}\subseteq T_{1}\subseteq\cdots\subseteq T_{n}=[n]$. Thus,
$\hat{f}(\boldsymbol{x})=\lambda_{0}f(\chi^{T_{0}})+\lambda_{1}f(\chi^{T_{1}})+\cdots+\lambda_{n}f(\chi^{T_{n}})$
is a well-defined extension of the function $f$ (called the Lovász extension of $f$) on the continuous domain $[0,1]^{n}$. The Lovász extension serves as a bridge between supermodularity and concavity: $f$ is supermodular if and only if its Lovász extension $\hat{f}$ is concave (Lovász 1983). Moreover, when the Lovász extension $\hat{f}$ is concave, it is expressible as the pointwise minimum of affine functions (Tawarmalani et al. 2013).
When the history-dependent effects are negative, the attraction value function $\alpha_{i}(\cdot)$ is supermodular, and its concave envelope is given by its Lovász extension (Tawarmalani et al. 2013). The key idea behind the closed-form description of the Lovász extension is to interpolate the attraction value function affinely over the nested historical assortments $\chi^{T_{0}},\chi^{T_{1}},\ldots,\chi^{T_{M}}$, where $T_{0}\subseteq T_{1}\subseteq\cdots\subseteq T_{M}$. These nested assortments can be thought of as a path through the set of all possible past assortments $\{0,1\}^{M}$ going from $(0,0,\ldots,0)$ to $(1,1,\ldots,1)$, with each step along one coordinate direction. For example, when the memory length is two, the nested assortments $(0,0),(1,0)$, and $(1,1)$ correspond to the blue path in Figure 1(a). This path starts at $(0,0)$, moves along the first coordinate, then the second, and finally reaches $(1,1)$. To treat the general case of mixed effects, we modify the nested assortments—or equivalently the path—as follows: Instead of moving from $(0,0,\ldots,0)$ to $(1,1,\ldots,1)$, we start at $\chi^{I}$ and end at $\chi^{M\setminus I}$, where $I$ denotes the indices of positive effects. For instance, consider mixed effects with $\beta_{i}^{1}>0$ and $\beta_{i}^{2}\leq 0$. The blue path in Figure 1(b), starting from $(1,0)$ and ending at $(0,1)$, corresponds to the past assortments that we will use in this case.
Figure 1: Illustration of switched nested assortments under two different history-dependent effects
(a) Satiation effects: $\beta_{i}^{1}\leq 0,\beta_{i}^{2}\leq 0$
$(0,0)$$(1,0)$$(1,1)$$(0,1)$$x_{i}^{t-1}$$x_{i}^{t-2}$
(b) Mixed effects: $\beta_{i}^{1}>0,\beta_{i}^{2}\leq 0$
$(0,0)$$(1,0)$$(0,1)$$(0,1)$$1-x_{i}^{t-1}$$x_{i}^{t-2}$
Now, we formally define the historical assortments used in our construction. For each product $i\in[N]$, let $I_{i}$ denote the indices of positive effects, that is, $I_{i}:=\{m\mid\beta_{i}^{m}>0\}$. Let $\Omega$ denote all permutations of past periods $[M]$. For each product $i\in[N]$ and a permutation $\sigma\in\Omega$, we define $M+1$ historical assortments, denoted as $\boldsymbol{h}^{\sigma}_{i,0},\boldsymbol{h}^{\sigma}_{i,1},\ldots,\boldsymbol{h}^{\sigma}_{i,M}$, as follows: $\boldsymbol{h}^{\sigma}_{i,0}=\chi^{I_{i}}$ and for each $k\in[M]$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{h}^{\sigma}_{i,k}=\begin{cases}\boldsymbol{h}^{\sigma}_{i,k-1}+\boldsymbol{e}_{\sigma(k)}&\text{ if }\sigma(k)\notin I_{i}\\ \boldsymbol{h}^{\sigma}_{i,k-1}-\boldsymbol{e}_{\sigma(k)}&\text{ if }\sigma(k)\in I_{i},\end{cases}}$ (SwitchNested)
With this definition, we are ready to describe the concave envelope of $\alpha_{i}(\cdot)$ which, after perspectification, yields an explicit linear description for the constraint (6).
Proposition 3.1
Constraint (6) is equivalent to the following system of linear inequalities.
$y^{t}_{i}\leq\alpha_{i}(\boldsymbol{h}^{\sigma}_{i,0})(\gamma^{t}_{i}-\tilde{z}_{i\sigma(1)}^{t})+\sum_{k\in[M]}\alpha_{i}(\boldsymbol{h}^{\sigma}_{i,k})(\tilde{z}_{i\sigma(k)}^{t}-\tilde{z}_{i\sigma(k+1)}^{t})\qquad\text{ for }\sigma\in\Omega\$ (7)
where $\tilde{z}_{i\sigma(M+1)}^{t}=0$, $\tilde{z}^{t}_{i\sigma(k)}=z^{t}_{i\sigma(k)}$ if $\sigma(k)\notin I_{i}$, and $\tilde{z}^{t}_{i\sigma(k)}=\gamma_{i}^{t}-z^{t}_{i\sigma(k)}$ if $\sigma(k)\in I_{i}$.
3.1.2 The Final Formulation
Using the linear constraints (3), (4), and (7) and the exponential cone constraint (5) to represent (Choice), we obtain an MIECP formulation of problem (HAP) as follows:
$\max_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{\rho},\boldsymbol{\gamma},\boldsymbol{z}}$ $\frac{1}{T}\sum_{t\in[T]}\sum_{i\in[N]}r_{i}y_{i}^{t}$ (Conic)
$\operatorname{s.t.}$ $\boldsymbol{x}^{t}\in\{0,1\}^{N}\cap\mathcal{X}\text{ and }\boldsymbol{x}\in\mathcal{P}$ $\operatorname{for}\,t\in[T]$
$\rho^{t}+\sum_{i\in[N]}y_{i}^{t}=1\quad$ $\operatorname{for}\,t\in[T]$
[…]
Theorem 3.2
(Conic) is a mixed-integer exponential cone formulation of problem (HAP).
Remark 3.3
In this remark, we consider the case where the memory length is less than two. In this case, we can derive the tightest convex extension, also known as the convex envelope, of $\alpha_{i}(\cdot)$. Using the convex envelope instead of the natural continuous extension, we obtain the following tighter formulation than the one given by constraint (5) for $M=2$. If $\boldsymbol{\beta}_{i}$ are all positive or negative, we have
$y_{i}^{t}\geq$ $\alpha_{i}(\boldsymbol{0})(\gamma_{i}^{t}-z_{i1}^{t}-z_{i2}^{t})+\alpha_{i}(\boldsymbol{e}_{1})z_{i1}^{t}+\alpha_{i}(\boldsymbol{e}_{2})z_{i2}^{t}$ (8a)
$y_{i}^{t}\geq$ $\alpha_{i}(\boldsymbol{e}_{1})(\gamma_{i}^{t}-z_{i2}^{t})+\alpha_{i}(\boldsymbol{e}_{2})(\gamma_{i}^{t}-z_{i1}^{t})+\alpha_{i}(\boldsymbol{1})(z_{i1}^{t}+z_{i2}^{t}-\gamma_{i}^{t}),$ (8b)
and if $\boldsymbol{\beta}_{i}$ are mixed, the convex envelope is
$y_{i}^{t}\geq$ $\alpha_{i}(\boldsymbol{0})(\gamma_{i}^{t}-z_{i1}^{t})+\alpha_{i}(\boldsymbol{e}_{1})(z_{i1}^{t}-z_{i2}^{t})+\alpha_{i}(\boldsymbol{1})z_{i2}^{t}$ (9a)
$y_{i}^{t}\geq$ $\alpha_{i}(\boldsymbol{0})(\gamma_{i}^{t}-z_{i2}^{t})+\alpha_{i}(\boldsymbol{e}_{2})(z_{i2}^{t}-z_{i1}^{t})+\alpha_{i}(\boldsymbol{1})z_{i1}^{t}$ (9b)
[…]
We conclude this section with a discussion of how our formulation relates to existing formulations. Recently, a large amount of research has focused on developing strong formulations for binary linear fractional programming problems and MNL-based assortment planning problems (e.g. Sen et al. 2018, Mehmanchi et al. 2019, Atamtürk and Gómez 2020, Kılınç-Karzan et al. 2023, He et al. 2024, Chen et al. 2024). To apply these results in solving problem (HAP), we need to linearize the nonlinear functions appearing in denominators and numerators—specifically, the attraction value functions. A prevalent linearization technique replaces each attraction value function by its multilinear extension, called the Fourier Expansion in O’Donnell (2014), and then linearizes the resulting multilinear functions using the recursive McCormick relaxation (McCormick 1976, Khajavirad 2023). We combine this linearization strategy with binary linear fractional programming formulation techniques to obtain an exact formulation of problem (HAP), which we refer to as the multilinear-extension-based formulation. The complete derivation is provided in the ecompanion EC.5.1. Later, in Section 6.1, a computational experiment shows that our formulation (Conic) significantly outperforms the multilinear-extension-based formulation.
[…]
$\max_{\boldsymbol{a}\in\mathcal{A}\subseteq\{0,1\}^{n}}\quad\sum_{k=1}^{K}\lambda_{k}\cdot\frac{\exp(u_{k}(\boldsymbol{a}))}{1+\exp(u_{k}(\boldsymbol{a}))},$
[…]
3.2 Optimality of Revenue-Ordered Policies
A prevalent strategy for solving assortment optimization is to construct Revenue-Ordered (RO) assortments, which is an assortment including products with revenue higher than a threshold. Specifically, RO structures and their variants have been shown to be optimal for solving assortment optimization problems under the MNL choice model (Talluri and Van Ryzin 2004), sequential MNL choice models (Gao et al. 2021), MNL choice models with network effects (Wang and Wang 2017), etc. Inspired by the simplicity and optimality of RO structures, we would like to understand how it works in our problem.
We find that when there are no cross-product and cross-period constraints, if the history-dependent effect is non-negative, that is, $\boldsymbol{\beta}\geq 0$, the optimal assortment in each period is an RO assortment, and the revenue threshold increases when $T$ increases. Such an assortment planning can be obtained by a sequential revenue-ordered policy. The sequential RO policy is obtained by sequentially selecting the best RO assortment based on the currently updated product utility and maximizing revenue in the current period. This policy is adaptive while ignoring the impact of the current assortment decision on future revenue. Details of the sequential RO policy are shown in Algorithm 1. Because there are no cross-product and cross-period constraints, the computation time required to find the optimal RO assortment is $\mathcal{O}(N)$ in each period (Talluri and Van Ryzin 2004, Liu and Van Ryzin 2008). Thus, the total computation time of the sequential RO policy is $\mathcal{O}(NT)$.
Algorithm 1 Sequential revenue-ordered policy of (HAP)
1:  Input: $\boldsymbol{r},\{\beta_{i}^{0}\}_{i\in[N]},\boldsymbol{\beta},M,N,T$
2:  for $t=1,2,\dots,T$ do
3:     Update history assortments: $\boldsymbol{x}^{t-1},\dots,\boldsymbol{x}^{t-M}$, and set $\boldsymbol{x}^{t-m}=\boldsymbol{0}$ if $t-m\leq 0$.
4:     Compute the best revenue-ordered assortment $\boldsymbol{x}^{t}=\mathop{\mathrm{argmax}}\limits_{\boldsymbol{z}\in\{0,1\}^{N}}\sum_{i\in[N]}r_{i}\pi^{t}_{i}(\boldsymbol{z},\boldsymbol{x}^{t-1},\ldots,\boldsymbol{x}^{t-M})$
5:  end for
6:  Output: $(\boldsymbol{x}^{1},\dots,\boldsymbol{x}^{T})$
We are now in a position to state the main result of this subsection, Theorem 3.4, characterizing conditions on which the sequential RO policy is optimal.
Theorem 3.4
Assume the absence of cross-product and cross-period constraints. Then, the sequential-revenue-ordered policy solves (HAP) if the history-dependent effects are non-negative, that is, $\boldsymbol{\beta}\geq\boldsymbol{0}$.
For the case of negative history-dependent effects, which is the computationally intractable regime of our model, as we argued in Proposition 2.1, the sequential policy fails to solve our model. However, the RO structure still exists under the negative and even mixed history-dependent effects and appears in a different form. For an optimal assortment planning of model (HAP), we argue that the set of products that are offered across all periods has an RO structure.
Proposition 3.5
Assume the absence of cross-product and cross-period constraints. Then, there exists an optimal assortment $\boldsymbol{x}=(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{T})$ of (HAP) such that $\bigl{\{}i\in[N]\bigm{|}\sum_{t\in T}x^{t}_{i}\geq 1\bigr{\}}$ is revenue-ordered.
4 Cyclic Policies: Optimality and Computation
In this section, we consider the situation in which a firm plans assortments for a long period. We establish the asymptotic optimality of cyclic policies in Section 4.1 and formulate the cyclic policy problem as an MIECP in Section 4.2. In Section 4.3, we characterize conditions under which the length of optimal cyclic policies is the memory length plus one. In Section 4.4, we show that such conditions lead to a tighter MILP formulation for finding optimal $(M+1)$-cyclic policies.
4.1 Optimality of Cyclic Policies
The assortment planning problem with an infinite horizon can be formulated as follows. The firm can select any assortment sequence (or policy) $\boldsymbol{x}=\{\boldsymbol{x}^{t}\}_{t\in\operatorname{\mathbb{Z}}_{+}}$ with elements in $\{0,1\}^{N}$. As before, the firm may face cross-product and cross-period constraints. Let $\mathcal{X}$ (resp. $\mathcal{P}^{\infty}$) denote the feasible set of cross-product (resp. cross-period) constraints. Then, the firm aims to find an optimal assortment sequence for the following problem:
$\sup_{\boldsymbol{x}}\biggl{\{}\lim_{T\to\infty}\frac{1}{T}\sum_{t\in[T]}\sum_{i\in[N]}r_{i}\pi_{i}^{t}(\boldsymbol{x}^{t},\dots,\boldsymbol{x}^{t-M})\biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\;\operatorname{for}t\in\operatorname{\mathbb{Z}}_{+},\ \boldsymbol{x}\in\mathcal{P}^{\infty}\biggr{\}}.$ (Infty)
An assortment policy $\boldsymbol{x}=\{\boldsymbol{x}^{t}\}_{t\in\operatorname{\mathbb{Z}}_{+}}$ is cyclic if there exists a positive integer $L$ such that $\boldsymbol{x}^{t+L}=\boldsymbol{x}^{t}$ for all $t\in\operatorname{\mathbb{Z}}_{+}$. The smallest $L$ for which this holds is called the cycle length of $\boldsymbol{x}$. A cyclic policy $\boldsymbol{x}$ can be represented by a finite sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$, where $L$ is the cycle length of $\boldsymbol{x}$. In contrast, whenever we say that a finite sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$ is cyclic, we are referring to the policy in which this finite sequence of assortments is repeated infinitely often.
For a given cycle of length $L$, we are interested in finding a cyclic policy of length $L$ that maximizes the long-term average revenue. This is equivalent to searching for a sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$ of length $L$ that maximizes the average revenue by repeating the sequence. Such a sequence can be defined as an optimal solution of a variant of (HAP). To define the variant, we modify the attraction value function so that it depends on past assortments in the cycle. Given a position $t$ in the cycle of length $L$, we use a function $\tau(\cdot|t)$ to track positions within the memory length $M$:
$\tau(m|t)=\begin{cases}{(t-m\text{ mod }L)}&\text{ if }(t-m\text{mod }L)>0\\ {L}&\text{ if }(t-m\text{ mod }L)=0.\end{cases}$
The attraction value of product $i$ at position $t$ can be specified as follows:
$\alpha_{i}^{t}(x^{1}_{i},\ldots,x^{L}_{i})=\exp\biggl{(}\beta_{i}^{0}+\sum_{m\in[M]}\beta^{m}_{i}x^{\tau(m\mid t)}_{i}\biggr{)}.$
A cyclic policy is called an $L$-cyclic policy if it is optimal for the following variant of problem (HAP):
$\max\Biggl{\{}\frac{1}{L}\sum_{t\in[L]}\sum_{i\in[N]}\frac{r_{i}x^{t}_{i}\alpha^{t}_{i}(x^{1}_{i},\ldots,x^{L}_{i})}{1+\sum_{j\in[N]}x^{t}_{j}\alpha^{t}_{j}(x^{1}_{i},\ldots,x^{L}_{i})}\Biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\operatorname{for}t\in[L],\ \boldsymbol{x}\in\mathcal{P}^{L}\Biggr{\}},$ (Cycle)
where $\mathcal{X}$ denotes cross-product constraints and $\mathcal{P}^{L}$ denotes cross-period constraints. In addition, we assume that a finite sequence of assortments satisfying $\mathcal{P}^{L}$ can be repeated infinitely often to obtain a cyclic policy that satisfies the cross-period constraint $\mathcal{P}^{\infty}$ in (Infty).
The main result of this subsection relates to problems (Infty) and (Cycle). To establish this result, we represent (Infty) as a directed graph, which we refer to as the assortment graph. Each node of the graph denotes a list of $M$ assortments, indexed from the latest to the earliest assortments in memory. Each arc represents a memory transition driven by an assortment decision. Specifically, an arc starting at a node should end at a node that drops the earliest assortment of the starting node and adds a new assortment. Hence, each arc encodes an assortment decision in a period, and its weight is defined as the single-period revenue generated by the new assortment. With such a graph, we can show that (Infty) is equivalent to finding the maximum mean cycle of the assortment graph (Karp 1978), whose cycle length is denoted as $L^{*}$. Given $L^{*}$, we can use (Cycle) to compute the cyclic policy. We formalize this observation as the following theorem.
Theorem 4.1
Given an instance of (Infty), there exists a positive integer $L^{*}$ such that $L^{*}$-cyclic policy is optimal to (Infty).
4.2 Finding Optimal Cyclic Policies
Formulation techniques developed for solving (HAP) can be used to obtain a mixed-integer exponential cone formulation for (Cycle). Since the main discrepancy between the two models is the definition of attraction value functions, we only need to modify constraints in (Conic) that involve attraction value functions and historical assortments. In particular, we replace $x^{t-m}_{i}$ in (4) by $x_{i}^{\tau(m|t)}$ and $\alpha_{i}(\cdot)$ in (7) by $\alpha^{t}_{i}(\cdot)$:
$\max\bigl{\{}0,\ \rho^{t}_{U}x^{\tau(m|t)}_{i}+\gamma^{t}_{i}-\rho^{t}_{U}\bigr{\}}\leq z_{im}^{t}\leq\min\bigl{\{}\gamma_{i}^{t},\rho^{t}_{U}x^{\tau(m|t)}_{i}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M]$ (10a)
$y^{t}_{i}\leq\alpha_{i}^{t}(\boldsymbol{w}^{\sigma}_{i,0})(\gamma^{t}_{i}-\tilde{z}_{i\sigma(1)}^{t})+\sum_{k\in[M]}\alpha_{i}^{t}(\boldsymbol{w}^{\sigma}_{i,k})(\tilde{z}_{i\sigma(k)}^{t}-\tilde{z}_{i\sigma(k+1)}^{t})$ $\quad\operatorname{for}\,\sigma\in\Omega,$ (10b)
where $\boldsymbol{w}^{\sigma}_{i,k}\in\{0,1\}^{L}$, and its $\tau(m|t)^{\text{th}}$ element equals the $m^{\text{th}}$ element of $\boldsymbol{h}_{i,k}^{\sigma}$ defined in (7) and others are zero. This replacement yields the following mixed-integer exponential cone formulation of (Cycle):
$\max_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{\rho},\boldsymbol{\gamma},\boldsymbol{\boldsymbol{z}}}$ $\frac{1}{L}\sum_{t\in[L]}\sum_{i\in[N]}r_{i}y_{i}^{t}$ (Cycle-Conic)
$\operatorname{s.t.}$ $\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\text{ and }\boldsymbol{x}\in\mathcal{P}^{L}$ $\operatorname{for}t\in[L]$
$\rho^{t}+\sum_{i\in[N]}y_{i}^{t}=1$ $\operatorname{for}\,t\in[L]$
[…]
(Cycle-Conic) has its own value because it is an efficient way to generate a feasible assortment planning with a large planning horizon. Especially, the retailer may pre-determine a cycle length $L$ based on experience or industry knowledge and use (Cycle-Conic) to derive the optimal policy under a given $L$.
4.3 Cycle Length Characterization
It is intuitive that, in order to avoid the negative impact on a product’s utility if the product exhibits a satiation effect, a product should not be offered consecutively. In particular, we say that a cross-period constraint on a policy $\boldsymbol{x}$ is non-overlapping if $M+1$ adjacent assortments $(\boldsymbol{x}^{t},\boldsymbol{x}^{t+1},\dots,\boldsymbol{x}^{t+M})$ are non-overlapping, that is, for each product $i\in[N]$, $x^{t}_{i}+x^{t+1}_{i}+\dots+x^{t+M}_{i}\leq 1$ for every $t\in\operatorname{\mathbb{Z}}_{+}$. This non-overlapping condition is common in assortment optimization literature (Liu et al. 2020, Chen et al. 2023b). For instance, Li et al. (2024) identify a similar non-overlapping condition on customer orders to plan assortments for a local warehouse. Under the non-overlapping condition, any cyclic policy with length $L\leq M$ must have empty sets in some periods. We do not consider such cyclic policies. For cyclic policies with length $L>M$, we find that this condition allows us to relate the length of an optimal cyclic policy for solving (Infty) to the customers’ memory length as follows.
Theorem 4.2
Assume that $M<N$. An $(M+1)$-cyclic policy is optimal for problem (Infty) if the cross-period constraint is non-overlapping.
We remark that Theorem 4.2 holds for mixed satiation-addiction effects, because its proof replies on a graph whose structure unaffected by signs of $\boldsymbol{\beta}$.
The condition $M<N$ guarantees that there at least exists a non-overlapping assortment policy without empty assortments. Without the non-overlapping condition, the optimal cycle length varies as the parameters change. In particular, a slight fluctuation in the input parameters, like the base utility of a product, could change the optimal cycle length. We provide an example to illustrate this situation in EC.4.4.
The base utility and history-dependent effect are estimated from real data and lie in confidence intervals. Hence, given the fact that the optimal cycle length is sensitive to input parameters, finding a cycle length with robust performance may be more critical than characterizing the optimal cycle length in practice. A positive result is that although the $(M+1)$-cyclic policy is not necessarily optimal for (Infty) without the non-overlapping constraint, it still provides a good approximation, as illustrated in the following example.
Example 4.3
We numerically test the performance of the $(M+1)$-cyclic policy under satiation effects. We fix the number of products $N=30$ and vary the planning horizon $T\in\{10,30,50\}$. The revenue and the base utility of each product are randomly generated from uniform distributions with ranges $[1,10]$ and $[-1,1]$, respectively. For a weak (W) (resp. strong (S)) satiation effect, we set the history-dependent effect $\beta_{i}^{m}$ using a uniform distribution $U[-1,0]$ (resp. $U[-2,-1]$). For each instance, we compute the revenue gap between the $(M+1)$-cyclic policy and the best feasible solution, defined as $100\%\times\frac{R_{\mathsf{fea}}-R_{\mathsf{cycle}}}{R_{\mathsf{fea}}}$, where $R_{\mathsf{cycle}}$ is the revenue achieved by the $(M+1)$-cyclic policy and $R_{\mathsf{fea}}$ is the revenue of the best feasible solution obtained by (Conic) within $7200$ seconds.
Table 1: Revenue gap of $(M+1)$-cyclic policies under satiation effects with $N=30$ and no constraints
[…]
W 10 0.51 1.69 S 10 0.39 4.45
W 30 0.16 0.60 S 30 -0.06 -0.50
W 50 0.08 0.52 S 50 -0.52 0.46
[…]
4.4 A Strong Formulation for Finding Non-overlapping $(M+1)$-Cyclic Policies
In this subsection, we propose a bound-free formulation for the $(M+1)$-cyclic policy, which is provably tighter than its conic counterpart (Cycle-Conic) whose computational performance depends on the bounds ($\rho^{t}_{L}$ and $\rho^{t}_{U}$) of the no-purchase probability variable $\rho^{t}$.
From the non-overlapping constraint $x_{i}^{t}+x^{\tau(1\mid t)}_{i}+\dots+x^{\tau(M\mid t)}_{i}\leq 1$, we have:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{i}^{t}\alpha_{i}^{t}(x^{1}_{i},\ldots,x^{L}_{i})=x_{i}^{t}\exp\biggl{(}\beta_{i}^{0}+\sum_{m\in[M]}\beta^{m}_{i}x^{\tau(m\mid t)}_{i}\biggr{)}=x_{i}^{t}\exp(\beta_{i}^{0})\qquad\text{ for }t\in[M+1].}$
Let $u_{i}=\exp(\beta_{i}^{0})$.  (Cycle) is equivalent to
$\max$ $\frac{1}{M+1}\sum_{t\in[M+1]}\sum_{i\in[N]}\frac{r_{i}u_{i}x^{t}_{i}}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}$ ((M+1)-Cyclic)
$\operatorname{s.t.}$ $x_{i}^{t}+\sum_{m\in[M]}x^{\tau(m\mid t)}_{i}\leq 1$ $\operatorname{for}i\in[N]$
$\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}$ $\operatorname{for}t\in[M+1].$
Note that this reformulation reveals a connection between (Cycle) and the assortment problem studied in Caro et al. (2014) in which a retailer must decide, in advance, the release date of each product in a given collection over a selling season. More specifically, when $M=1$, ((M+1)-Cyclic) is exactly the case where a retailer faces two planning horizons and sells products with a single period life-cycle, and, in each period, customers make a purchase decision on offered products according to the MNL model.
Next, we introduce variables for building a strong formulation of ((M+1)-Cyclic). For each period $t\in[M+1]$, we use $\rho^{t}$ and $u_{i}\cdot\gamma^{t}_{i}$ to represent the no-purchase probability and choice probability of product $i$, respectively,
$\rho^{t}=\frac{1}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}\quad\text{ and }\quad\gamma_{i}^{t}=\frac{x^{t}_{i}}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}\qquad\text{ for }i\in[N].$ (CC)
Note that this is known as the Charnes-Cooper transformation (Charnes and Cooper 1962). For each period $t\in[M+1]$, it maps all possible assortments $\boldsymbol{x}^{t}\in\{0,1\}^{N}$ into all possible choice probabilities $(\rho^{t},u_{1}\gamma^{t}_{1},\ldots,u_{N}\gamma^{t}_{N})$, where $(\rho^{t},\boldsymbol{\gamma}^{t})$ belongs to
$\rho^{t}+\sum_{j\in[N]}u_{j}\gamma^{t}_{j}=1\quad\text{ and }\quad\gamma^{t}_{j}\in\{0,\rho^{t}\}\,\operatorname{for}j\in[N].$ (11)
Using these relations, we are ready to present a new formulation of ((M+1)-Cyclic):
$\max\limits_{\boldsymbol{x},\boldsymbol{\rho},\boldsymbol{\gamma},\boldsymbol{\Gamma}}\quad$ $\frac{1}{M+1}\sum_{t\in[M+1]}\sum_{i\in[N]}r_{i}u_{i}\gamma_{i}^{t}$
$\operatorname{s.t.}\quad$ $x_{i}^{t}+\sum_{m\in[M]}x^{\tau(m\mid t)}_{i}\leq 1\text{ and }\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}$ $\operatorname{for}i\in[N]\text{ and }t\in[M+1]$ (12a)
$\rho^{t}+\sum_{i\in[N]}u_{i}\gamma^{t}_{i}=1\text{ and }0\leq\gamma^{t}_{i}\leq\rho^{t}\quad$ $\operatorname{for}i\in[N]\text{ and }t\in[M+1]$ (12b)
$x^{t}_{i}=\gamma^{t}_{i}+\sum_{j\in[N]}u_{j}\Gamma^{t}_{ij}$ $\operatorname{for}i\in[N]\text{ and }t\in[M+1]$ (12c)
$\max\{0,\gamma^{t}_{i}+\gamma^{t}_{j}-\rho^{t}\}\leq\Gamma^{t}_{ij}\leq\min\{\gamma^{t}_{i},\gamma^{t}_{j}\}\quad$ $\operatorname{for}i\in[N]\,,j\in[N]\setminus\{i\},\text{ and }t\in[M+1]$ (12d)
$\Gamma^{t}_{ii}=\gamma^{t}_{i}$ $\operatorname{for}i\in[N]\text{ and }t\in[M+1].$ (12e)
The constraint (12b) is obtained from relaxing the constraints from (11). The constraint (12c) is derived from the following equality, with the variable $\gamma^{t}_{i}$ replacing the first fractional term and a new variable $\Gamma^{t}_{ij}$ replacing each fractional term in the summation,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x^{t}_{i}=\frac{x^{t}_{i}}{1+\sum_{k\in[N]}u_{k}x^{t}_{k}}+\sum_{j\in[N]}u_{j}\frac{x_{i}^{t}x_{j}^{t}}{1+\sum_{k\in[N]}u_{k}x^{t}_{k}}\qquad\operatorname{for}\,t\in[M+1]}.$
To obtain (12d), we use the no-purchase variable $\rho^{t}$ to scale the following McCormick envelope of $x_{i}^{t}x_{j}^{t}$ (McCormick 1976):
$\max\{0,x^{t}_{i}+x^{t}_{j}-1\}\leq x^{t}_{i}x^{t}_{j}\leq\min\{x_{i}^{t},x_{j}^{t}\}\qquad\operatorname{for}i\in[N],j\in[N]\setminus\{i\}.$
The last constraint (12e) is obtained by observing that the relation $x^{t}_{i}x^{t}_{i}=x^{t}_{i}$ holds for binary variables. Unlike the formulation (Cycle-Conic), this formulation does not rely on bounds of the no-purchase probability variables. More specifically, in the derivation of (Cycle-Conic), we use $\rho^{t}x^{t}_{i}$ as a nonlinear scaling factor, demanding a linearization using bounds of $\rho^{t}$. In the development of linear inequalities in (12), we avoid using a nonlinear scaling factor and, instead, directly use $\rho^{t}$ as a scaling variable to derive (12d).
The following result shows this bound-free formulation is theoretically tighter than the formulation (Cycle-Conic). One of the most important properties of an mixed integer program (MIP) formulation is the strength of its natural continuous relaxation that is obtained by ignoring the integrality constraints. This is important because a tighter continuous relaxation often indicates a faster convergence of the branch-and-bound algorithm on which most commercial MIP solvers are built (Vielma 2015). Since (HAP) has a maximization objective, a tighter formulation has a smaller continuous relaxation objective value. Let $Z_{\textsc{Bound-free}}$ (resp. $Z_{\textsc{Cycle-Conic}}$) be the optimal objective value of the natural continuous relaxation of formulation (12) (resp. (Cycle-Conic)). The two objective values have the following relationship.
Theorem 4.4
$Z_{\textsc{Bound-free}}\leq Z_{\textsc{Cycle-Conic}}$ if the cross-period constraint is non-overlapping.
5 Case Study
This section presents a case study using transaction data from a corporation cafeteria. We first estimate the history-dependent effect and then use the results to design assortments. Our estimations indicate that customers have a significant but short-term satiation effect. Furthermore, our findings suggest that including history-dependent effects in choice models will lead to greater assortment varieties and higher revenue.
5.1 Data Description and Estimation
We cooperated with a Chinese company that provides catering services to over $4,000$ corporation cafeterias and were given access to a dataset for one cafeteria that mainly serves workday luncheons for customers who work nearby. The dataset includes $27,144$ transactions for $53$ main dishes, covering $110$ workdays from July to December 2021. The data for each transaction includes the time, dish, and quantity, as well as a card ID if the dish was purchased with a membership card. The dataset also includes information about the price, main ingredient, and flavor of each dish. In what follows, we use the terms dish and product and also the terms menu and assortment interchangeably.
Our dataset indicates that 84.8% of the transactions came from 418 members, which generated 84.5% of the total revenue. Thus, we focus on the members’ transactions because they are frequent visitors and generate most of the revenue. Our dataset also reveals that the probability of a member dining in the cafeteria more than once for two, three, or four consecutive days was 72.5%, 87.7%, and 91.2%, respectively. This suggests that customers’ dish selections could be influenced by past menus.
We predict customers’ choices using the MNL model with history-dependent effects. The utility function for product $i$ on day $t$ is
$U_{i}^{t}=\boldsymbol{\theta}^{T}X_{i}+\theta_{r}r_{i}+\sum_{m\in[M]}x_{i}^{t-m}\beta^{m}+\xi_{i}^{t},$
where $X_{i}$ denotes covariates that include a fixed effect and the main ingredient and flavor of product $i$; $r_{i}\in[0,1]$ is a normalized price; $\boldsymbol{\theta}$ and $\theta_{r}$ are coefficients of $X_{i}$ and price, respectively; $\boldsymbol{\theta}^{T}X_{i}+\theta_{r}r_{i}$ is the base utility $\beta_{i}^{0}$ defined in Section 2; and $\xi_{i}^{t}$s are i.i.d. standard Gumbel random variables. The history-dependent effect is captured by the term $\sum_{m\in[M]}x_{i}^{t-m}\beta^{m}$, in which $M$ is the memory length. Our model is a standard MNL model if $M=0$. Note that we set the same history-dependent coefficient for all products because not all products have consecutive offering records, and therefore, we cannot estimate the product-wise history-dependent effect.
We utilize members’ transactional data to calibrate the MNL choice model with the utility function presented above. We split the transaction data into two sets: data from July to October for training and data from November to December for test. We first use five-fold cross-validation on the first data set to estimate the utility function, and then we optimize weekly menus based on the estimators. Next, we conduct the five-fold cross-validation again, this time on the second data set, to select the value of $M$ with the best out-of-sample performance, and we set this as the “ground truth” model for evaluation, which is a common approach in policy evaluation (Cao et al. 2022). To address the fact that our dataset lacks no-purchase records, we employ the expectation-maximization (EM) approach, a widely used technique to deal with censored data (van Ryzin and Vulcano 2017, Şimşek and Topaloglu 2018), to calibrate our model.
Table 5.1 summarizes the estimation and five-fold validation results for $M\in\{0,1,2,3,4\}$, using the first data set. Column $\theta_{r}$ presents price coefficient estimators, and columns $\beta^{1}$ to $\beta^{4}$ give history-dependent effects. The columns labeled $\chi^{2}$ and Log-likelihood provide the average out-of-sample value in the five-fold cross-validation (normalized by sample size). The percentage values in parentheses indicate the relative improvement of each model with a history-dependent effect compared to the standard MNL model.
[…]
Estimators and cross-validation results under different memory length Estimators Out-of-sample Performance $M$ $\theta_{r}$ $\beta^{1}$ $\beta^{2}$ $\beta^{3}$ $\beta^{4}$ $\chi^{2}$ Log-likelihood 0 $-2.558^{***}$ 0.0860 -1.7796 (0.0767) 1 $-2.550^{***}$ $-0.475^{***}$ 0.0828 ($3.67\%$) -1.7769 ($0.15\%$) (0.0767) (0.0498) 2 $-2.520^{***}$ $-0.493^{***}$ $-0.366^{***}$ 0.0822 ($4.38\%$) -1.7754 ($0.24\%$) (0.0767) (0.0499) (0.0568) 3 $-2.540^{***}$ $-0.492^{***}$ $-0.358^{***}$ $-0.289^{***}$ 0.0819 ($4.71\%$) -1.7749 ($0.27\%$) (0.0768) (0.0500) (0.0568) (0.0708) 4 $-2.540^{***}$ $-0.495^{***}$ $-0.361^{***}$ $-0.286^{***}$ $-0.146^{***}$ 0.0826 ($3.91\%$) -1.7753 ($0.25\%$) (0.0768) (0.0500) (0.0569) (0.0709) (0.0647) Note: The numbers in parentheses under the estimators report the standard errors of the estimations. “-” indicates that a model does not include the corresponding predictor. ${***}p<0.001$. The columns “$\chi^{2}$” and “Log-likelihood” show the average value of the five-fold cross-validation (normalized by sample size) followed by a percentage in parentheses, which is the relative improvement of each model with a history-dependent effect compared to the standard MNL model.
Table 5.1 shows significant negative history-dependent effects, and the scale of the effects decreases as the time interval between the current and historical offerings increases. The estimation results are consistent for different values of $M$. For instance, $\beta^{1}$ is consistently about $-0.49$ for all $M$. This suggests that the attraction value of a product will shrink by about $\exp(-0.49)=61\%$ if it was offered yesterday. In other words, if a product is sold on two consecutive days, we need to lower its price by at least RMB $(0.492/2.540)\times 15\approx 2.9$ when $M=3$ to compensate for the negative history-dependent effect, where we normalize prices by the maximum price RMB $15$ in the estimation. These results demonstrate the importance of considering history-dependent effects in assortment planning.
The average out-of-sample validation results are reported in the last two columns of Table 5.1. They suggest that history-dependent effects can improve prediction power (i.e., lower $\chi^{2}$ and higher log-likelihoods). When $M=3$, the model presents the lowest $\chi^{2}$ and the highest log-likelihood. This result suggests that a small memory length ($M=3$) can capture the history-dependent effect in customers’ choice.
5.2 Assortment Planning
In this section, we use the formulation proposed in Section 3.1 to optimize weekly menus using the above estimators. We design weekly workday menus because the cafeteria does not frequently operate on weekends, and in practice, the cafeteria indeed plans for a week. We use the estimated parameters in Table 5.1 as the input for our conic models. To keep the menu sizes comparable, we impose a daily cardinality constraint of $6$, which is the median assortment size in the dataset.
5.2.1 Assortment Patterns
Figure 5.2.1 shows the accumulated number of products for different offering times, the prices, and the base attraction values obtained from the assortments under the MNL model, under history-dependent MNL models with different memory lengths, and under an assortment planning used in practice. The current practice assortment was randomly chosen from the second test data set.
[…]
(a) Offering times
(b) Price
(c) Base attraction value
Information of products offered in assortments obtained by different models
As shown in Figure 2(a) and 2(b), the MNL-based model tends to offer fewer products with higher prices. Moreover, the number of products that are offered increases with the memory length $M$, because it is beneficial to offer more products and avoid repeating the same product within the memory length when satiation effects last for a long period. When the memory length increases, the assortments include more products with lower prices and higher base attraction values, as shown in Figure 2(b) and 2(c).
[…]
Assortment details of four different policies (a)-(d) show the assortments of four polices. The horizontal axes represent products indexed. The vertical axes represent the days of a week. The products offered are colored blue. (e) and (f) show the price and base attraction value of each product.
Figure 5.2.1 presents assortment details for examining how the memory length affects assortment patterns. Figure 5.2.1 (a)-(d) show the assortments obtained from MNL-based models with $M\in\{0,1,2,3\}$. The horizontal axes represent products indexed from the highest to the lowest price. The vertical axes represent the days of a week. The products offered are colored blue. The subfigures show that the interval between two adjacent offerings of the same dish grows when $M$ increases, and optimal assortments represent a cyclic-style structure. The interval for two adjacent offerings of most products is the memory length $M$, which suggests that the cycle length is $M+1$, aligning with the $(M+1)$-cyclic policy in Section 4. Moreover, we observe that products with higher prices or lower base attraction values are more sensitive to the satiation effect and thus will be offered at most once within the span of the memory length. For instance, the high-price products 1 to 5 strictly repeat once in intervals of $M$ days to avoid the satiation effect, possibly because the revenue loss caused by the utility deterioration will be scaled up by the high price. On the other hand, products with relatively high base attraction value may possibly be offered more than once within the memory period, e.g., product 17,18. The price and base attraction value for each product are shown in Figure 5.2.1 (e) and (f) for reference.
5.2.2 Performance Comparison
In this section, we focus on two metrics to evaluate the assortment performance: variety and revenue, as these are concerns for the corporation cafeteria. We adopt the Herfindahl-Hirschman Index (HHI), a measure of market concentration (Rhoades 1993), to quantify the variety of an assortment planning. The HHI value is $\sum_{i\in[n]}(\frac{h_{i}}{H})^{2}$, where $n$ is the total number of products, $h_{i}$ is the number of times product $i$ is offered, and $H=\sum_{i\in[n]}h_{i}$. A low HHI value reflects a large variety. To compute the out-of-sample revenue, we randomly sample 20 subsets from the second data set. In each subset, we estimate the “ground truth” model with $M=2$ that was selected by cross-validation on the second data set and calculate the revenue based on the “ground truth” demand model. The average revenue across the 20 data sets is used as the out-of-sample revenue to avoid random noise. Besides the assortments discussed in Section 5.2.1, we additionally compute assortments obtained by the $(M+1)$-cyclic policy with non-overlapping constraints that is presented in Section 4.
[…]
Assortment variety and average revenue in the out-of-sample test The horizontal axis represents the Herfindahl-Hirschman Index value of each policy. A small HHI value means a wide variety. The vertical axis shows the average revenue of a policy across 20 test sets. We normalize the revenue of the current practice to one.
Figure 5.2.2 illustrates the relationship between the variety and the revenue of an assortment plan. The horizontal axis represents the HHI value, and the vertical axis shows the average normalized revenue. We normalize the out-of-sample revenue of different assortment plans with the revenue of current practice as a benchmark. Generally, assortments with a low HHI value can help avoid the negative impacts of satiation effects and maintain high product utility. Figure 5.2.2 shows that the assortments generated by our model ($M>0$) have a lower HHI and higher revenue than the assortments obtained with the MNL model. For instance, the average revenue obtained using our model with $M=3$ is approximately $10.4\%$ greater than the revenue under the MNL model. The $(M+1)$-cyclic policy also performs well, with cyclic assortments experiencing only a slight revenue loss compared to the optimal assortments. We also note that the manually decided assortments in practice have low HHI values but carry substantial revenue losses, as in practice, managers may decide the assortments based on many factors that are not observed. These results demonstrate that a wide variety of products with a low frequency can mitigate the satiation effect. Our model achieves this by efficiently balancing products’ utility, revenue, and history-dependent effects.
6 Numerical Study
This section presents our numerical studies on synthetic data. Section 6.1 illustrates the efficiency of our formulations regarding computation time and optimality gap compared to other formulations. Section 6.2 shows that our formulation generates significantly more revenue than benchmark heuristics. In Section 6.3, we demonstrate that the bound-free model (12) can scale up to solve large-size instances. Our numerical studies were implemented in the Julia (Bezanson et al. 2017), using modeling language JuMP (Lubin et al. 2023), on a Virtual Machine with a 32-core Intel Xeon (Skylake, IBRS) @2.39 GHz processor. All MILP models were solved using Gurobi 11.0.0 (Gurobi Optimization, LLC 2023), and MIECP models were solved using Mosek 10.2 (MOSEK 2024).
6.1 Formulation Efficiency
We consider four alternate formulations for solving (HAP) without constraints and set the bounds of the no-purchase probability variable as $\rho^{t}_{L}=\frac{1}{({1+\sum_{j\in[N]}\exp(\beta_{j}^{0}+\sum_{m\in I_{j}}\beta_{j}^{m}))}}$ and $\rho^{t}_{U}=1$ for the first three models:
  • $\mathsf{Conic}$: This is the mixed-integer exponential cone formulation given by (Conic).
  • $\mathsf{Env}$: This is the mixed-integer linear program obtained by replacing the exponential cone constraint in (Conic) with the convex envelope based constraints (8) and (9), as detailed in Remark 3.3.
  • $\mathsf{ML}$: This is the mixed-integer linear formulation based on a multilinear extension of attraction value functions, and its derivation is detailed in the ecompanion EC.5.1.
  • $\mathsf{SCIP}$: This refers to solving the binary nonlinear model (HAP) by the global optimization solver SCIP.
We consider the satiation effects in this numerical study. For a fixed number of products $N$ and memory length $M$, the base utility of each product $\beta_{i}^{0}$ is drawn from a uniform $[-1,1]$ distribution, and its revenue $r_{i}$ from a uniform $[1,10]$ distribution. The no-purchase utility is 00. For each $i\in[N]$ and $m\in[M]$, we set the history-dependent effect $\beta_{i}^{m}$ from two uniform distributions $U[-1,0]$ and $U[-2,-1]$ to represent a weak and strong satiation effect, denoted as W and S, respectively. We vary $N\in\{30,50,70\}$, $M\in\{1,2\}$, and $T\in\{5,10\}$. For each combination of $N$, $M$, and a satiation level (W or S), we generate five instances. In EC.6, we provide additional computational results on settings where history-dependent effects are mixed, memory lengths are large, and both cross-periods and cross-products constraints are present. To solve our MIECP models with large memory lengths, we propose a cutting-plane algorithm, which is discussed in EC.5.3.
We solve each instance of (HAP) using one of the formulations $\mathsf{Conic}$, $\mathsf{Env}$, $\mathsf{ML}$, and $\mathsf{SCIP}$. We report the computation time and the end gap, which is defined as $G_{\textsf{end}}=100\%\times\frac{R_{\mathsf{U}}-R_{\mathsf{IP}}}{R_{\mathsf{IP}}}$, where $R_{\mathsf{U}}$ and $R_{\mathsf{IP}}$ is the best upper bound and the best integer solution at termination with a time limit of 3600 seconds and an optimality gap tolerance of $0.5\%$, respectively.
Table 2: Computation time and end gap of four formulations under satiation effects
Conic Env ML SCIP
T M N $\boldsymbol{\beta}$$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$
5 1 30 W 5 36 0 5 6 0 5 267 0 0 - 648
5 1 30 S 5 90 0 5 96 0 5 527 0 0 - 920
5 1 50 W 3 65 0.64 5 657 0 3 473 1.04 0 - 1380
5 1 50 S 3 1766 0.66 4 1249 0.68 1 3492 1.51 0 - 1527
5 1 70 W 3 453 1.75 4 102 2.17 3 855 2.26 0 - 1798
5 1 70 S 1 1695 0.87 2 757 0.68 1 2523 1.87 0 - 2191
5 2 30 W 5 267 0 5 95 0 3 1339 0.99 - - 913
5 2 30 S 3 527 2.39 3 363 1.5 0 - 11.76 0 - 1046
5 2 50 W 3 689 0.95 5 160 0 1 2760 1.57 0 - 1415
5 2 50 S 0 - 2.35 1 2133 1.64 0 - 49.13 0 - 1756
5 2 70 W 0 - 1.81 3 811 1.6 0 - 16.35 0 - 2369
5 2 70 S 1 1020 3.41 1 57 2.37 0 - 79.22 0 - 2459
10 1 30 W 0 - 0.92 4 422 0.86 0 - 0.94 0 - 744
10 1 30 S 0 - 2.14 1 2712 1.77 0 - 2.86 0 - 1005
10 1 50 W 0 - 1.35 2 1994 1.4 0 - 1.48 0 - 1477
10 1 50 S 0 - 3.82 0 - 3.12 0 - 3.88 0 - 1577
10 1 70 W 0 - 2.44 1 313 2.06 0 - 3.27 0 - 1907
10 1 70 S 0 - 3.08 0 - 1.9 0 - 4.43 0 - 2305
10 2 30 W 0 - 2.37 0 - 1.4 0 - 4.15 0 - 1001
10 2 30 S 0 - 7.52 0 - 5.8 0 - 165.07 0 - 1151
10 2 50 W 0 - 2.9 0 - 2 0 - 63.85 0 - 1563
10 2 50 S 0 - 8.73 0 - 6.05 0 - 608.16 0 - 1891
10 2 70 W 0 - 4.66 0 - 3.05 0 - 283.46 0 - 2756
10 2 70 S 0 - 6.51 0 - 4.64 0 - 750.15 0 - 2763
[…]
T M N $\boldsymbol{\beta}$ Conic Env ML SCIP T M N $\boldsymbol{\beta}$ Conic Env ML SCIP
5 1 30 W 3.86 3.22 156.73 - 10 1 30 W 4.38 3.65 179.23 -
5 1 30 S 6.81 5.89 301.08 - 10 1 30 S 7.83 6.56 343.77 -
5 1 50 W 2.89 2.5 295.99 - 10 1 50 W 3.3 2.82 336.43 -
5 1 50 S 5.64 5.04 517.19 - 10 1 50 S 7.11 5.64 588.26 -
5 1 70 W 3.95 3.55 400.53 - 10 1 70 W 5.17 4.77 455.25 -
5 1 70 S 3.84 3.45 713.79 - 10 1 70 S 5.05 4.2 810.67 -
5 2 30 W 4.57 4.57 224.85 - 10 2 30 W 5.64 5.61 275.29 -
5 2 30 S 13.17 13.17 587.57 - 10 2 30 S 15.4 15.32 786.57 -
5 2 50 W 4.77 4.77 430.75 - 10 2 50 W 5.9 5.76 539.85 -
5 2 50 S 10.1 10.06 921.6 - 10 2 50 S 13.34 12.09 1419.53 -
5 2 70 W 5.47 5.3 717.98 - 10 2 70 W 6.94 6.46 940.42 -
5 2 70 S 6.85 6.43 1243.68 - 10 2 70 S 9.02 7.85 1890.09 -
Besides the computation time and end gap, it is also helpful to compare the continuous relaxation of different formulations. We compute the root gap by $100\%\times\frac{R_{\textsf{Rlx}}-R_{\textsf{IP}}}{R_{\textsf{IP}}}$, where $R_{\mathsf{Rlx}}$ is the objective value of the continuous relaxation of a given formulation. A small root gap indicates a tight formulation. Table 3 reports the average root gap of Conic, Env, and ML under different parameter configurations. Overall, the root gap of our formulations Conic and Env is significantly smaller than that of ML. It indicates the tightness of our formulations. Moreover, the root gap of our formulations is stable and does not scale with the problem size. For instance, the root gap of Env is no more than $7\%$ when $M=1$ and no more than $16\%$ when $M=2$.
6.2 Performance of Heuristic Policies
We compare the revenue generated by our formulation against two heuristic policies without constraints.
  • $\mathsf{Sequential-RO}$: The sequential RO-policy shown in Algorithm 1. It can serve as a heuristic policy for settings with satiation effects.
  • $\mathsf{Sequential-LOSPO}$: For each period, this policy drops one satiation product that causes the greatest revenue reduction from the base assortment obtained from the $\mathsf{Sequential-RO}$ policy. We refer to it as sequential leave-one-satiation-product-out (Sequential-LOSPO) policy.
We consider instances with mixed addiction-satiation effects. A product has an addiction (resp. satiation) effect with probability $\theta$ (resp. $1-\theta$), where $\theta\in\{0,0.1,0.2\}$. When $\theta=0$, all products have satiation effects. For a product with addiction (resp. satiation) effect, its $\{\beta_{i}^{m}\}_{m\in[M]}$ are uniformly sampled from $[0,1]$ (resp. $[-2,-1]$). We fix $M=2$ and generate five instances for each combination of $N,T$, and $\theta$.
We compare revenue obtained by the two heuristic policies and our formulation. We define the relative revenue gap as $100\%\times\frac{R_{\mathsf{Env}}-R_{\mathsf{heuristic}}}{R_{\mathsf{Env}}}$, where $R_{\mathsf{heuristic}}$ is revenue generated by one of two heuristic policies. $R_{\mathsf{Env}}$ is computed by $\mathsf{Env}$ within 3600 seconds. Table 4 shows that the revenue gap of both heuristics is not negligible, especially when $\theta$ is small. For instance, the smallest relative gap is at least more than 10% if $\theta=0$, suggesting that our formulation has more advantage in scenarios where the satiation effect is dominant. Although the revenue gap shrinks when $\theta$ increases, our formulation still holds value as it has an optimality guarantee and is easily extended to constrained cases.
[…]
$\mathsf{Sequential-RO}$ $\mathsf{Sequential-LOSPO}$
N T $\theta=0$ $\theta=0.1$ $\theta=0.2$ $\theta=0$ $\theta=0.1$ $\theta=0.2$
30 5 18.60 13.02 2.99 15.09 11.77 2.60
30 10 31.37 19.43 4.18 28.47 18.65 3.81
50 5 14.39 5.18 2.02 12.90 4.63 1.84
50 10 25.72 7.73 2.82 24.20 7.27 2.66
70 5 12.31 4.00 3.59 10.91 3.80 3.40
70 10 21.86 5.74 5.03 20.61 5.59 4.88
6.3 Computing $(M+1)$-Cyclic Policies
In this section, we demonstrate the efficiency and scalability of the bound-free model (12) for computing $(M+1)$-cyclic policies under the non-overlapping constraint. Note that due to the presence of $\boldsymbol{\Gamma}$ variables, the number of variables in (12) is $\mathcal{O}(N^{2})$. To address this issue, we propose a projected-cutting-plane algorithm to obtain a projected approximation of (12). The cutting plane algorithm first sets a base formulation without using $\boldsymbol{\Gamma}$ variables, that is, constraints (12c)-(12e), and solves its continuous relaxation to obtain a continuous solution. Then, the algorithm calls a separation oracle to generate cuts violated by the continuous solution, which are selected from the projection of constraints (12c)-(12e) onto the space of the base formulation, i.e., the space of $(\boldsymbol{x},\boldsymbol{\rho},\boldsymbol{\gamma})$ variables. Adding new cuts to the base model leads to a tighter formulation. Running the separation oracle once is a round of cut generation. We can introduce new cuts over $K$ rounds and refer to the resulting formulation as BF-$K$, where “BF” stands for “bound-free”. We refer to EC.5.2 for a detailed description of the projected cutting-plane algorithm.
Clearly, a larger $K$ leads to a tighter formulation at the cost of adding cuts. To select a proper $K$ that balances the formulation tightness and the formulation size, we compute the optimality gap closed after the $K^{\text{th}}$ round as follows
$\textsf{GClosed}_{K}=100\%\times\frac{\textsf{Opt}_{0}-\textsf{Opt}_{K}}{\textsf{Opt}_{0}-\textsf{Opt}_{\infty}},$
where $\textsf{Opt}_{K}$ is the optimal objective value of the continuous relaxation of BF-$K$, and the index 00 (resp. $\infty$) refers to the base (resp. limiting) model. Table 5(a) reports the gap closed by BF-$K$ for $K\in\{1,2,3\}$ on instances with $M\in\{1,2\}$ and $N\in\{50,100\}$. BF-$1$ closes about $80\%$ of the optimality gap, suggesting setting $K=1$ is good enough.
Table 5: The computational performance of (12) with a projected-cutting-plane implementation
(a) Performance of the cutting-plane algorithm
N M $\textsf{GClosed}_{1}$ $\textsf{GClosed}_{2}$ $\textsf{GClosed}_{3}$
50 1 79.25 91.19 96.21
50 2 82.24 93.14 95.90
100 1 89.82 95.61 97.58
100 2 77.24 90.30 93.48
(b) Performance of bound-free model BF-$1$
N M #$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$
50 1 5 0.57 0.30
50 2 5 70.74 0.46
100 1 5 8.19 0.42
100 2 5 378.59 0.49
Table 5(b) reports the performance of BF-$1$. The result demonstrates the scalability and efficiency of the bound-free model compared to (Conic). BF-1 solves all instances to optimality within 400 seconds. In contrast, under the number of products $N=50$ and memory length $M=2$, both Conic and Env fail to solve some of the instances as shown in Table 2.
7 Conclusion
This paper studies a multi-period assortment planning problem with history-dependent customer utility. We prove that the problem is NP-hard under negative history-dependent effects. To address this for general history-dependent effects, we develop an MIECP formulation to find a global optimal solution. We also prove the optimality of the sequential policy for the addiction case and the asymptotic optimality of cyclic policies while relating the cycle length to customer memory length.
Our research can be extended in several directions. First, it is important to develop efficient approximation algorithms, such as fully polynomial-time approximation schemes, to address the history-dependent assortment planning problem, especially in cases with large memory lengths and long planning horizons. Second, extending the model to online settings for personalized assortments or dynamic settings to adjust assortments to quickly respond to stochastic demand. Third, investigating joint assortment, pricing, or inventory problems under history-dependent effects. Finally, exploring history-dependent effects in other choice models for assortment optimization.